### v1

# MambaFlow: A Mamba-Centric Architecture for End-to-End Optical Flow Estimation

Abstract
Optical flow estimation based on deep learning, particularly the recently proposed top-performing methods that incorporate the Transformer, has demonstrated impressive performance, due to the Transformer’s powerful global modeling capabilities. However, the quadratic computational complexity of attention mechanism in the Transformers results in time-consuming training and inference. To alleviate these issues, we propose a novel MambaFlow framework that leverages the high accuracy and efficiency of Mamba architecture to capture features with local correlation while preserving its global information, achieving remarkable performance. To the best of our knowledge, the proposed method is the first Mamba-centric architecture for end-to-end optical flow estimation. It comprises two primary contributed components, both of which are Mamba-centric: a feature enhancement Mamba (FEM) module designed to optimize feature representation quality and a flow propagation Mamba (FPM) module engineered to address occlusion issues by facilitate effective flow information dissemination. Extensive experiments demonstrate that our approach achieves state-of-the-art results, despite encountering occluded regions. On the Sintel benchmark, MambaFlow achieves an EPE all of 1.60, surpassing the leading 1.74 of GMFlow. Additionally, MambaFlow significantly improves inference speed with a runtime of 0.113 seconds, making it 18% faster than GMFlow. The source code will be made publicly available upon acceptance of the paper.
Figure 1: End-point-error on Sintel (clean) vs. inference time (ms) and model params (M). All models were trained on FlyingChairs (C), FlyingThings3D (T), Sintel (S), KITTI (K), and HD1K (H) datasets, with testing performed using one NVIDIA GeForce RTX 3090 GPU. Compared to state-of-the-art models, our proposed model achieves comparable accuracy while significantly reducing inference time, demonstrating an optimal balance between computational efficiency and performance.
1 Introduction
Optical flow estimation, a fundamental problem in computer vision, aims to compute the motion vectors of pixels between consecutive frames in a video sequence. Optical flow estimation holds significant and widespread applications in areas such as action recognition [42, 37], video interpolation [11, 34, 19], and autonomous driving [38].
Architectures based on Convolutional Neural Networks (CNNs) have achieved pioneering progress in the field of optical flow estimation. FlowNet [4], as the first end-to-end CNNs-based flow network, revealed the immense potential of this learning paradigm. However, it faced limitations in handling large displacements and complex motions, along with high computational costs. Subsequently, PWC-Net [30] improved estimation accuracy in complex motion scenarios by incorporating pyramid feature extraction and a cost-volume construction, which effectively captured multi-scale local features. Nevertheless, it still encountered challenges in large displacement scenarios. To address these issues further, RAFT [31] introduced a recurrent iterative refinement mechanism, significantly enhancing the accuracy of optical flow estimation. Still, its optimization process relies on numerous iterations, leading to increased parameter counts and substantially longer inference time.
In recent years, the Transformer architecture, with its powerful global modeling capabilities, has been introduced into optical flow estimation tasks. For instance, GMFlow [36] and FlowFormer [10] leverage attention mechanisms to capture long-range dependencies, significantly improving the accuracy of optical flow estimation. However, the quadratic computational complexity of the Transformers significantly increases the inference time in high-resolution scenarios, severely limiting their efficiency in practical applications. This raises a critical question: How can we achieve efficient optical flow estimation with reduced computational overhead while preserving high accuracy?
Mamba [5], rooted in state space models (SSMs)[6], employs an efficient selective state space mechanism to excel in tasks requiring the integration of global information and handling complex spatiotemporal dependencies. Simultaneously, it achieves linear computational complexity when capturing long-range dependencies in sequences. Inspired, we propose a novel optical flow estimation framework, MambaFlow, designed to address the challenges of large displacements, occlusions, and computational efficiency. As demonstrated in Fig. 1, our method significantly reduces inference time while maintaining high accuracy. The framework introduces two core innovations: a feature enhancement Mamba (FEM) module, which optimizes feature representations to enhance global matching performance, and a flow propagation Mamba (FPM) module, which propagates flow information to tackle challenges in occluded regions.
Specifically, our framework begins with a CNN backbone to extract robust features from the input image pair. We then propose the FEM module, which employs the Self-Mamba and Cross-Mamba mechanisms to strengthen the correlations between the extracted features, allowing for the capture of complex global information and long-range dependencies. Building on the enhanced features provided, a global matching module compute similarities and perform global optimization, to generate initial flow. This process fully utilizes global information, effectively tackling the challenges posed by large displacements. At last, we propose a Mamba-centric flow propagation method, the FPM module, which leverages feature similarity to guide propagation and produce the final accurate flow [36], tackling the occlusion issue in predicting optical flow.
Our main contributions can be summarized as follows:
  • We propose MambaFlow, a novel optical flow estimation architecture built on the Mamba framework. To the best of our knowledge, this is the first Mamba-centric architecture designed for end-to-end optical flow estimation. Our method achieves comparable accuracy to state-of-the-art algorithms while significantly reducing inference time. With a runtime of 0.113 seconds, our method is 18% faster than GMFlow and 10 times faster than MS-RAFT+.
  • We introduce a feature enhancement Mamba (FEM) module, which integrates Self-Mamba and Cross-Mamba mechanisms to address insufficient feature representation. This module established a new Mamba-centric paradigm for accurate and efficient feature enhancement.
  • We design a flow propagation Mamba (FPM) module for flow propagation that utilizes feature self-similarity to guide the final flow estimation. This module tackles the poor performance in recurrent refined methods by effectively propagating flow information, offering significant potential to enhance accuracy and speed, both currently and in the foreseeable future.
2 Related Work
2.1 Optical Flow Estimation
The optical flow problem is traditionally approached as an energy minimization task. Based on the assumptions of brightness constancy and spatial smoothness, Horn and Schunck [8] introduced a variational method to tackle this problem. The following studies refined optical flow estimation by framing it as an optimization problem, improving accuracy and robustness through regularization that maximizes the visual similarity between corresponding pixels in the image. In the era of deep learning, following the introduction of FlowNet [4] as the first CNN-based optical flow method, numerous subsequent algorithms such as PWC-Net [30] and SpyNet [28], adopted a coarse-to-fine strategy for optical flow estimation. However, small displacements of fast movement were often missed via the inaccurate flow guidance at the coarse level. To overcome this obstacle, RAFT [31] proposed an iterative refinement method using full-field, achieving significant improvements. Constructing upon the innovation of RAFT, later works such as GMA [15] improved the handling of occluded regions by introducing a global motion aggregation module. Similarly, SeparableFlow [40] introduced a separable cost volume module that used a global motion aggregation layer to utilize global contextual cues and prior knowledge to resolve motion ambiguity in regions. Despite the advancements, CNN-based models, including those built upon the RAFT framework, still rely on multiple iterations of refinement to achieve high-precision optical flow estimation. This iterative process significantly increases computational overhead and slows inference speed, thereby limiting their practicality for efficiency-sensitive applications.
In recent years, the Transformer architecture has demonstrated exceptional proficiency in modeling long-range dependencies and global information through its self-attention mechanism, making it a popular choice for optical flow estimation tasks. By incorporating more global and contextual information, it has significantly improved estimation accuracy. FlowFormer [10] further optimized model performance by encoding 4D cost tags into cost memory and applying alternating groups of Transformer layers in the latent space. CRAFT [29] strengthened the globality and semantic stability of features by leveraging semantic smoothing Transformer layers. GAFlow [22] added Gaussian attention to the existing Transformer block to emphasize local attributes during representation learning, highlight local neighborhoods containing fine-grained structural information, and strengthen motion affinity during matching. In addition, many models took advantage of the Transformer’s global modeling capabilities, adopting a global matching strategy to avoid multiple iterations of refinement, significantly enhancing efficiency. GMFlow [36] reformulated the optical flow problem as a global matching problem and obtained the optical flow using one-time global matching, which was more accurate and faster. Transflow [21] used spatial self-attention and cross-attention mechanisms between adjacent frames to effectively captured global dependencies, providing more accurate correlations and reliable matches in flow estimation. Still, despite these advances in accuracy, the Transformer-based optical flow methods face challenges due to the quadratic complexity of the attention mechanism led to huge computational complexity and memory requirements. This slows inference speed and restricts their use in resource-limited settings, creating practical bottlenecks.
In our proposed MambaFlow architecture, we employ the Mamba module, which contains linear complexity and discriminative feature representation. This approach significantly reduces the overhead of computational resources and memory requirements while maintaining high accuracy comparable to Transformer-based methods, A remarkable breakthrough for the practical application of optical flow models.
2.2 State Space Model in Optical Flow
State space models [7], which originated from control theory, have been effectively integrated with deep learning to capture long-range dependencies. Mamba [5] introduced an input adaptation mechanism to enhance the state space model, providing faster inference speed, higher throughput and overall indicators compared to the Transformers of the same size, especially in natural language. Related work on Mamba [33] has shown that it has comparable performance to the Transformer-based methods while significantly reducing computational complexity. This superiority led to further exploration like Vision Mamba [44] and Vmamba [20], which applied SSMs to visual tasks such as image classification [20, 44], video understanding [17, 32], biomedical image segmentation [23, 35] and others [9, 12, 26, 27, 39]. However, their applicability to optical flow estimation remained unclear until recently. Lin et al. [18] pioneered Mamba into the scene flow estimation task, proposing an SSM-based iterative update module to replace the GRU module as a decoder. This approach can effectively integrate global motion information in the point cloud and significantly improved inference speed.
Inspired by the Mamba’s linear complexity, we propose a Mamba-centric optical flow estimation architecture, aiming to replace the traditional attention mechanism and thus improve computational efficiency of the model. In this work, we explore the potential of Mamba for modeling sequential interactions in optical flow tasks and propose FEM module and FPM module. The FEM module enhances the features from pairs of images, capturing the dependency between them, while the FPM module refines optical flow predictions for occluded pixels by leveraging both the optical flow field and image structure similarity. By utilizing Mamba’s linear time complexity, our approach overcomes the quadratic complexity limitations of traditional Transformer models when processing long sequences, resulting in significantly improved inference efficiency.
Figure 2: The overall network architecture of our MambaFlow. First, we perform downsampling using convolutional layers and input the initially extracted features into the feature enhancement Mamba group (FEMG) module, which is composed of multiple feature enhancement Mamba (FEM) module. Each FEM module consists of Self-Mamba, Cross-Mamba, and MLP for feature enhancement. Next, global matching is applied to the output feature matrix for similarity matching, and the original optical flow is obtained through a softmax layer. Finally, the flow propagation Mamba (FPM) module is used to address inaccuracies in optical flow prediction caused by occlusion.
3 Method
3.1 Preliminaries
SSMs are inspired by linear systems, mapping an input sequence $x_{t}\in\mathbb{R}$ to an output $y_{t}\in\mathbb{R}$ through a hidden state $h_{t}\in\mathbb{R}^{N}$. The model can be formulated as linear ordinary differential equations (ODEs):
$h_{t}=\overline{\mathbf{A}}h_{t-1}+\overline{\mathbf{B}}x_{t},$ (1)
$y_{t}={\mathbf{C}}h_{t},$ (2)
where $\overline{\mathbf{A}}$ and $\overline{\mathbf{B}}$ represent the discretized parameters derived from the given parameters $\mathbf{A}\in\mathbb{R}^{N\times N}$, $\mathbf{B}\in\mathbb{R}^{N\times 1}$, and the time scale parameter $\boldsymbol{\Delta}$:
$\overline{\mathbf{A}}=\exp(\boldsymbol{\Delta}\mathbf{A}),$ (3)
$\overline{\mathbf{B}}=(\boldsymbol{\Delta}\mathbf{A})^{-1}(\exp(\boldsymbol{\Delta}\mathbf{A})-\mathbf{I})\cdot\boldsymbol{\Delta}\mathbf{B}.$ (4)
Additionally, $\mathbf{C}\in\mathbb{R}^{1\times N}$ serves as the projection matrix mapping the hidden state $h_{t}$ to the output $y_{t}$.
Finally, the models generate the output using a global convolution operation, defined as:
$\overline{\mathbf{K}}=(\mathbf{C}\overline{\mathbf{B}},\mathbf{C}\overline{\mathbf{A}}\overline{\mathbf{B}},\ldots,\mathbf{C}\overline{\mathbf{A}}^{M-1}\overline{\mathbf{B}}),$ (5)
$y=x\ast\mathbf{\overline{K}},$ (6)
where $M$ represents the length of the input sequence $x$, and $\mathbf{\overline{K}}\in\mathbb{R}^{M}$ denotes a structured convolutional kernel.
The recently proposed advanced state-space model, Mamba, enhances $\mathbf{B}$, $\mathbf{C}$, and $\boldsymbol{\Delta}$ by making them input-dependent, thereby enabling dynamic feature representations. Specifically, these matrices are defined as follows:
$\mathbf{B}_{i}=S_{B}(x_{i}),$ (7)
$\mathbf{C}_{i}=S_{C}(x_{i}),$ (8)
$\boldsymbol{\Delta}_{i}=\text{Softplus}(S_{\Delta}(x_{i})),$ (9)
where $S_{B}$, $S_{C}$, and $S_{\Delta}$ are linear projection functions, and $i$ denotes the index of the $i$-th element in the sequence.
Simultaneously, the parallel scan algorithm enables Mamba to leverage the same parallel processing advantages as described in (5), thereby facilitating efficient training.
Figure 3: The architecture of feature enhancement Mamba (FEM) module. It consists of three core components: Self-Mamba (a), which focuses on a single feature input, Cross-Mamba (b), which handles the interaction between two features, and MLP (c), which performs non-linear feature transformation.
3.2 Overall Architecture
Optical flow can be intuitively understood as a task aimed at finding pixel-wise correspondences between images. As shown in Fig. 2, our MambaFlow consists of three components: feature enhancement Mamba, global matching and flow propagation Mamba.
Given two consecutive frames $\boldsymbol{I}_{1}$ and $\boldsymbol{I}_{2}$, we apply a shared-weight convolutional neural network to perform downsampling, obtaining dense feature maps $\boldsymbol{F}_{1},\boldsymbol{F}_{2}\in\mathbb{R}^{H\times W\times D}$. Here, $H,W,$ and $D$ denote the height, width, and feature dimension of the feature maps, respectively. Next, the image features are processed through the feature enhancement stage to obtain the enhanced features of the $l$-th layer, $\boldsymbol{\mathit{F}}_{q}$ and $\boldsymbol{\mathit{F}}_{v}\in\mathbb{R}^{H\times W\times D}$. This stage is composed of multiple stacked FEM modules, each consisting of Self-Mamba and Cross-Mamba blocks. Additionally, a Multi-Layer Perceptron (MLP) layer is introduced at the end of each group to further integrate local and global features and enhance feature representation. Finally, we employ the Global Matching method to predict optical flow and use our FPM module to significantly improve performance in occluded regions.
3.3 Self-Mamba
The input features $\boldsymbol{\mathit{F}}_{1},\boldsymbol{\mathit{F}}_{2}\in\mathbb{R}^{H\times W\times D}$ are enriched with learnable positional embeddings to retain spatial positional information. Subsequently, $\boldsymbol{\mathit{F}}_{1}$ and $\boldsymbol{\mathit{F}}_{2}$ are processed through a weight-sharing Self-Mamba block, generating $x_{q}$ and $x_{v}$, respectively.
As shown in Fig. 3 (a), the processing within each Self-Mamba block can be summarized as a bidirectional operation, where the forward and backward hidden states are computed and combined to enrich the feature representation. The overall process is expressed as:
[…]
Here, the Self-Mamba function represents the state space model used to compute the forward and backward hidden states, with the sequence $x$ being reversed to distinguish between the forward and backward directions.
3.4 Cross-Mamba
Mamba is specifically designed to capture long-term dependencies within a single sequence, but it cannot model the interrelations between two sets as effectively as the cross-attention mechanism in the Transformers. To overcome this limitation, we propose Cross-Mamba, an extension of the Mamba framework that adapts the cross-attention mechanism to better capture the interactions between two distinct sets. This idea, which builds upon the work by [33], is further extended in our approach. Since the parameters $\overline{\mathbf{A}}$, $\overline{\mathbf{B}}$, ${\mathbf{C}}$, and $\boldsymbol{\Delta}$ are associated with the input, $\overline{\mathbf{K}}$ in (5) is extended into an $M\times M$ matrix [1]. The element $\alpha_{i,j}$, located at row $i$ and column $j$, is determined as follows:
$\alpha_{i,j}=\mathbf{C_{i}}\prod_{k=j+1}^{i}\overline{\mathbf{A}}_{k}\overline{\mathbf{B}}_{j}.$ (11)
Substituting the equations from (3) to (9), we can ultimately derive the output of Mamba as follows:
$y_{i}=\sum_{j=1}^{i}\alpha_{i,j}x_{j}=\sum_{j=1}^{i}Q_{i}K_{j}H_{i,j}V_{j},$ (12)
where
$Q_{i}$ $=S_{C}(x_{i}),$ (13)
$\begin{split}K_{j}=&\left(\text{Softplus}(S_{\Delta}(x_{j}))\mathbf{A}\right)^{-1}\\ &\left(\exp\left(\text{Softplus}(S_{\Delta}(x_{j}))\mathbf{A}\right)-\mathbf{I}\right)\\ &\left(\text{Softplus}(S_{\Delta}(x_{j}))S_{B}(x_{j})\right),\end{split}$ (14)
$H_{i,j}$ $=\prod_{k=j+1}^{i}\exp\left(\text{Softplus}(S_{\Delta}(x_{k}))\mathbf{A}\right),$ (15)
$V_{i}$ $=x_{i}.$ (16)
Cross-attention is a crucial mechanism for modeling the relationships between two sets. It operates by computing interactions between the query ($Q$) from one set and the keys ($K$) and values ($V$) from another set, thereby identifying and focusing on the most relevant elements in the target set.By applying $x_{q}$ to (13), and $x_{v}$ to (14), (15) and (16), respectively, Cross-Mamba can be expressed in the following form:
$y_{i}=\sum_{j=1}^{i}f_{m,q}(x_{q,i})f_{m,k}(x_{v,j})H_{i,j}x_{v,j}.$ (17)
Cross-Mamba extends the original Mamba by expanding its application from a single input sequence to two input sequences, as shown in Fig. 3 (b). It is formally defined as follows:
$y=\text{Cross-Mamba}(x_{q},x_{v}).$ (18)
Similarly, we introduce the bidirectional mechanism of Self-Mamba into Cross-Mamba. The formulation is expressed as follows:
[…]
3.5 Global Matching
After extracting the enhanced features $\boldsymbol{\mathit{F}}_{q}$ and $\boldsymbol{\mathit{F}}_{v}$ from our FEM module, we construct a 4D cost volume of size $H\times W\times H\times W$ by computing the dot-product similarity between all pixel pairs from both features. The construction process is expressed as:
$\mathbf{C}=\frac{\boldsymbol{\mathit{F}}_{q}\,\boldsymbol{\mathit{F}}_{v}^{\top}}{\sqrt{D}},$ (20)
where $D$ is the feature dimension. The constructed 4D cost volume can be interpreted as a collection of $H\times W$ 2D cost maps, where each map evaluates the feature similarity between a single pixel in $\boldsymbol{\mathit{F}}_{q}$ and all pixels in $\boldsymbol{\mathit{F}}_{v}$. To obtain a matching distribution, a softmax operation is applied across the last two dimensions of the cost volume:
$\mathbf{M}=\text{softmax}(\mathbf{C}).$ (21)
To estimate the optical flow, the correspondence $\hat{\mathbf{G}}$ is obtained by taking a weighted average of the 2D pixel grid coordinates $\mathbf{G}\in\mathbb{R}^{H\times W\times 2}$ in the target image with the matching distribution $M$:
$\hat{\mathbf{G}}=\mathbf{M}\mathbf{G}.$ (22)
Finally, the optical flow $\mathbf{V}_{\text{original}}$ is calculated as the difference between the corresponding pixel coordinates:
$\mathbf{V}_{\text{original}}=\hat{\mathbf{G}}-\mathbf{G}.$ (23)
The global matching method effectively improves the accuracy and stability of optical flow estimation by capturing global contextual information.
3.6 Flow Propagation Mamba
To address the issue of occluded regions or pixels outside the image boundaries, we introduced a self-similarity-based module called the flow propagation Mamba (FPM) module (as illustrated in Fig. 4), which leverages feature information from the source image to capture pixel-level similarity relationships. Since adjacent pixels typically exhibit similar features, they demonstrate higher similarity, thus providing a more robust solution for optical flow estimation. Specifically, by inputting $\boldsymbol{\mathit{F}}_{q}$ into the Mamba block, we enhance the features in occluded regions and further extract the weights of these occluded areas. This helps optimize the final optical flow estimation, thereby improving accuracy and robustness in complex scenarios. The formula is as follows:
$\mathbf{V}_{\text{final}}=\text{softmax}\left(\text{Mamba}(\boldsymbol{\mathit{F}}_{q})\right)\mathbf{V}_{\text{initial}},$ (24)
where $\mathbf{V}_{\text{initial}}\in\mathbb{R}^{H\times W\times 2}$ represents the optical flow predicted by the global matching.
Figure 4: The architecture of flow propagation Mamba (FPM) module.
Figure 5: Visual results on Sintel test set.
Training Data Method Sintel KITTI Inference Cost
[…]
C+T+S+K+H RAFT [31] 1.61 2.86 5.10 5.3M 140.7ms
SeparableFlow [40] 1.50 2.67 4.64 6.0M 229.1ms
GMA [15] 1.39 2.47 5.15 5.9M 183.3ms
DIP [43] 1.44 2.83 4.21 5.4M 498.9ms
GMFlowNet [41] 1.39 2.65 4.79 9.3M 244.3ms
CRAFT [29] 1.45 2.42 7.79 6.3M 483.4ms
MS-RAFT [13] 1.37 2.67 4.88 13.5M 627.3ms
MatchFlow [3] 1.33 2.64 4.72 14.8M 283.8ms
MS-RAFT+ [14] 1.23 2.68 4.15 16.0M 1108.2ms
GMFlow [36] 1.74 2.90 9.32 4.7M 138.5ms
MambaFlow (ours) 1.60 2.78 8.53 8.9M 113.5ms
Table 1: MambaFlow achieves at least 18% faster inference on the KITTI (test) dataset compared to other state-of-the-art methods with competitive performance. All compared methods were evaluated under the same training configuration, and a unified ”C+T+S+K+H” training strategy was adopted to ensure the fairness of the results. For performance measurements, we used the normalized setting of batch size 1 and input resolution 540 × 960. In addition, all inference times are measured on the RTX 3090 graphics card. The significant performance improvement highlights MambaFlow’s advantages in speed and efficiency. It is worth noting that while maintaining competitive accuracy on the Sintel and KITTI datasets, we have successfully achieved a significant improvement in speed, further verifying the superiority of MambaFlow in the speed-accuracy balance.
4 Experiments
Datasets.
Following the previous methodology, we first pre-train on the FlyingChairs (Chairs)[4] and FlyingThings3D (Things)[24] datasets. For fine-tuning, we adopt a mixed training set consisting of KITTI [25], HD1K [16], FlyingThings3D, and Sintel [2] training sets. Finally, we further fine-tune the model on the KITTI training set and report the performance on the respective online benchmarks.
Metrics.
In our experiments, we employed evaluation metrics commonly used in optical flow estimation. Firstly, we utilized the end-point-error (EPE) to measure the average Euclidean distance between the predicted and ground truth optical flow vectors. For the KITTI dataset, we adopted the F1-all metric, which represents the percentage of pixels with an error exceeding 3 pixels or a relative error greater than 5%. To gain deeper insights into the model’s performance under large motion magnitudes, we defined $s_{40+}$ as the EPE for pixels where the ground truth flow magnitude exceeds 40 pixels.
setup Things (val) Sintel (train) Param (M)
[…]
full 6.45 - 2.23 - 3.41 - 8.1 -
w/o self 6.89 $\uparrow$6.82.30 $\uparrow$3.13.49 $\uparrow$2.36.9 $\downarrow$14.8%
w/o cross 11.10 $\uparrow$72.13.78 $\uparrow$69.55.16 $\uparrow$51.34.4 $\downarrow$45.7%
w/o MLP 7.13 $\uparrow$10.52.41 $\uparrow$8.13.66 $\uparrow$7.35.2 $\downarrow$35.8%
w/o pos 6.65 $\uparrow$3.12.29 $\uparrow$2.73.48 $\uparrow$2.18.1 -
Table 2: FEM components. Ablation study on individual components, revealing that Cross-Mamba contributes the most.
layers Sintel (train) KITTI (train) Param (M)
[…]
4 1.41 2.79 9.53 27.15 5.1
6 1.37 2.72 9.16 25.38 6.9
8 1.36 2.63 8.54 24.97 8.9
10 1.35 2.68 8.51 24.67 10.8
12 1.36 2.55 8.10 23.91 12.7
24 1.34 2.50 7.96 22.97 23.4
Table 3: Number of Mamba layers. Using 8 layers of Mamba strikes an optimal balance between accuracy and parameter efficiency.
Implementation Details.
Our model is constructed using the PyTorch library and trained on 4 NVIDIA GeForce RTX 3090 GPUs. Following the training strategy from previous works [36, 31], we use the AdamW optimizer. The model is first pre-trained on the Chairs dataset for 100k iterations with a batch size of 16 and a learning rate of $4\times 10^{-4}$. Then, it is fine-tuned on the Things dataset for 200k iterations with a batch size of 8 and a learning rate of $2\times 10^{-4}$. During this process, we select the best-performing model on the Things dataset and fine-tune it for an additional 800k iterations to validate the ablation study. Subsequently, the model fine-tuned for 200k iterations on the Things dataset is used as the initial model and further trained on the Sintel dataset, keeping the same hyperparameters as those used for the Things dataset, including a batch size of 8 and a learning rate of $2\times 10^{-4}$. Finally, we continue training on the KITTI-2015 training set for 200k iterations with the batch size and learning rate unchanged.
4.1 Comparison with State-of-the-Arts
We compared the performance of different optical flow algorithms on the Sintel and KITTI datasets, and the experimental results are shown in Table 1.
Results on Sintel.
MambaFlow achieved an EPE of 1.60 on the Sintel Clean dataset and an EPE of 2.78 on the Final dataset. Fig. 5 presents a qualitative analysis of optical flow estimation, comparing MambaFlow with other state-of-the-art approaches on the Sintel test set. The results show that MambaFlow’s performance on Clean and Final is at the same level as current mainstream methods such as RAFT and GMA, demonstrating its strong adaptability in both simple and complex scenarios.
Results on KITTI.
MambaFlow achieves an F1-all score of 8.53% on the KITTI dataset while demonstrating a significant advantage in inference speed. Specifically, MambaFlow requires only 113.5ms for inference on an RTX 3090, making it 18.1% faster than GMFlow (138.5ms), which is the fastest among the compared methods. Furthermore, MambaFlow runs 2.5 times faster than MatchFlow (283.8ms) and nearly 10 times faster than MS-RAFT+ (1108.2ms), while maintaining competitive accuracy. This demonstrates that MambaFlow is able to maintain comparable accuracy to mainstream methods while significantly reducing inference time when handling larger and more challenging scenarios.
4.2 Ablation Analysis
Components of FEM.
We ablate different Fearture Enhancement Mamba components in Table 2. Among them, Cross-Mamba contributes the most as it models the mutual relationship between two features, a relationship that is otherwise missing in the features extracted from the convolutional backbone. Self-Mamba aggregates contextual cues within the same feature, resulting in an additional performance improvement, and positional information introduces location dependence into the matching process, helping to mitigate ambiguities present in feature-similarity-based matching alone. Removing the Multi-Layer Perceptron (MLP) reduces the number of parameters, but also results in a slight decrease in performance.
Number of FEM layers.
We conduct an ablation study on the number of FEM layers, as shown in Table 3. The experiment compares the impact of varying the number of FEM layers on overall performance. Among all the tested configurations, the FEMG module with 8 Mamba layers exhibits the best performance.
Training iterations.
The experimental results in Table 4 show that with the increase in the number of training iterations, our framework has achieved certain improvements in performance. This shows that more training iterations can effectively improve the performance of the model, making it more stable and accurate in subsequent tasks.
training Things (val) Sintel (train) KITTI (train)
[…]
200K 3.04 1.19 2.60 9.02 25.92
400K 2.81 1.14 2.57 8.76 25.18
600K 2.73 1.10 2.52 8.14 24.18
800K 2.61 1.09 2.56 7.57 23.51
Table 4: Training iterations. MambaFlow benefits from more training iterations.
5 Conclusion
In this paper, we explore for the first time the application of the advanced Mamba in optical flow estimation, and then propose a simple, accurate and efficient Mamba-centric optical flow method, MambaFlow, for end-to-end flow prediction. It consists of two components built on Mamba architecture: the FEM module and FPM module. The FEM is proposed for accurate and efficient feature enhancement, incorporating the Self-Mamba, Cross-Mamba mechanisms to tackle the issue of inadequate feature representation. Furthermore, the FPM that employs feature self-similarity to direct the ultimate flow is crafted for the flow propagation, addressing the challenge of traditional recurrent refined methods. Our method achieves comparable accuracy to state-of-the-art algorithms while significantly reducing inference time, with a runtime of 0.113 seconds—18% faster than GMFlow and 10 times faster than MS-RAFT+. We attribute this to the high accuracy and efficiency of Mamba in capturing features with local correlation while maintaining its global information. As far as we know, this is the first Mamba-centric architecture designed for end-to-end optical flow estimation. We are assured that additional variants of the Mamba-centric methods may emerge , possessing significant potential to boost both precision and velocity in the future.

### v2

@@
 Abstract
-Optical flow estimation based on deep learning, particularly the recently proposed top-performing methods that incorporate the Transformer, has demonstrated impressive performance, due to the Transformer’s powerful global modeling capabilities.
-However, the quadratic computational complexity of attention mechanism in the Transformers results in time-consuming training and inference.
-To alleviate these issues, we propose a novel MambaFlow framework that leverages the high accuracy and efficiency of Mamba architecture to capture features with local correlation while preserving its global information, achieving remarkable performance.
-To the best of our knowledge, the proposed method is the first Mamba-centric architecture for end-to-end optical flow estimation.
-It comprises two primary contributed components, both of which are Mamba-centric: a feature enhancement Mamba (FEM) module designed to optimize feature representation quality and a flow propagation Mamba (FPM) module engineered to address occlusion issues by facilitate effective flow information dissemination.
-Extensive experiments demonstrate that our approach achieves state-of-the-art results, despite encountering occluded regions.
-On the Sintel benchmark, MambaFlow achieves an EPE all of 1.60, surpassing the leading 1.74 of GMFlow.
-Additionally, MambaFlow significantly improves inference speed with a runtime of 0.113 seconds, making it 18% faster than GMFlow.
+Recently, the Mamba architecture has demonstrated significant successes in various computer vision tasks, such as classification and segmentation.
+However, its application to optical flow estimation remains unexplored.
+In this paper, we introduce MambaFlow, a novel framework designed to leverage the high accuracy and efficiency of the Mamba architecture for capturing locally correlated features while preserving global information in end-to-end optical flow estimation.
+To our knowledge, MambaFlow is the first architecture centered around the Mamba design tailored specifically for optical flow estimation.
+It comprises two key components: (1) PolyMamba, which optimizes feature representation; and (2) PulseMamba, which facilitates efficient flow information dissemination.
+Our extensive experiments demonstrate that MambaFlow achieves remarkable results.
+On the Sintel benchmark, MambaFlow records an endpoint error (EPE) of 1.43 and an inference speed of 0.113 seconds, surpassing the state-of-the-art methods including GMFlow (with 18.9% lower EPE and 18.1% faster inference), SeparableFlow (5% lower EPE and 50.5% faster), CRAFT (1.11% lower EPE and 76.5% faster), and DIP (0.7% lower EPE and 77.2% faster)—demonstrating stronger potential for real-world deployment on resource-constrained devices.
 The source code will be made publicly available upon acceptance of the paper.
+1 Introduction
+Optical flow estimation, a fundamental problem in computer vision, aims to compute the motion vectors of pixels between consecutive frames in a video sequence.
+It holds significant and widespread applications in areas such as action recognition [65, 56], video interpolation [16, 51, 28], autonomous driving [59], etc.
+Architectures based on Convolutional Neural Networks (CNNs) have achieved pioneering progress in the field of optical flow.
+FlowNet [5], as the first end-to-end CNNs-based flow network, revealed the immense potential of this learning paradigm.
+However, it faced limitations in handling large displacements and complex motions.
+Subsequently, PWC-Net [44] improved estimation accuracy in complex motion scenarios by incorporating pyramid feature extraction and a cost-volume construction, which effectively captured multi-scale local features.
+Nevertheless, it still encountered challenges in large displacement scenarios.
+To address these issues, RAFT [45] introduced a recurrent iterative refinement mechanism, which significantly enhanced the accuracy of optical flow estimation.
+However, such innovation brought a new challenge, which its optimization process relies on numerous iterations, leading to increased parameter counts and longer inference time.
+Many subsequent variants of the RAFT algorithm have attempted upgrates, such as MS-RAFT [19] and SEA-RAFT [49], but their ability to handle large displacements and complex motions still requires further improvement.
+To tackle the aforementioned issues, the Transformer architecture, known for its powerful global modeling capabilities, has been introduced into optical flow estimation tasks [15, 42, 30, 43].
+Most of them leverage attention mechanism to enhance feature representation, which capture long-range dependencies to significantly improve the accuracy.
+Notably, unlike optical flow algorithms based on RNNs for flow decoding [45, 20], GMFlow [54] utilizes the attention mechanism to iteratively decode the final optical flow, resulting in faster flow refinement.
+However, they encounter the quadratic computational complexity of Transformers both during the training and inference stages [54, 43, 15, 42, 30], which severely limits their efficiency in practical applications.
 Figure 1: End-point-error on Sintel (clean) vs.
@@
 Compared to state-of-the-art models, our proposed model achieves comparable accuracy while significantly reducing inference time, demonstrating an optimal balance between computational efficiency and performance.
-1 Introduction
-Optical flow estimation, a fundamental problem in computer vision, aims to compute the motion vectors of pixels between consecutive frames in a video sequence.
-Optical flow estimation holds significant and widespread applications in areas such as action recognition [42, 37], video interpolation [11, 34, 19], and autonomous driving [38].
-Architectures based on Convolutional Neural Networks (CNNs) have achieved pioneering progress in the field of optical flow estimation.
-FlowNet [4], as the first end-to-end CNNs-based flow network, revealed the immense potential of this learning paradigm.
-However, it faced limitations in handling large displacements and complex motions, along with high computational costs.
-Subsequently, PWC-Net [30] improved estimation accuracy in complex motion scenarios by incorporating pyramid feature extraction and a cost-volume construction, which effectively captured multi-scale local features.
-Nevertheless, it still encountered challenges in large displacement scenarios.
-To address these issues further, RAFT [31] introduced a recurrent iterative refinement mechanism, significantly enhancing the accuracy of optical flow estimation.
-Still, its optimization process relies on numerous iterations, leading to increased parameter counts and substantially longer inference time.
-In recent years, the Transformer architecture, with its powerful global modeling capabilities, has been introduced into optical flow estimation tasks.
-For instance, GMFlow [36] and FlowFormer [10] leverage attention mechanisms to capture long-range dependencies, significantly improving the accuracy of optical flow estimation.
-However, the quadratic computational complexity of the Transformers significantly increases the inference time in high-resolution scenarios, severely limiting their efficiency in practical applications.
-This raises a critical question: How can we achieve efficient optical flow estimation with reduced computational overhead while preserving high accuracy?
-Mamba [5], rooted in state space models (SSMs)[6], employs an efficient selective state space mechanism to excel in tasks requiring the integration of global information and handling complex spatiotemporal dependencies.
-Simultaneously, it achieves linear computational complexity when capturing long-range dependencies in sequences.
-Inspired, we propose a novel optical flow estimation framework, MambaFlow, designed to address the challenges of large displacements, occlusions, and computational efficiency.
+Recently, Mamba [8, 10] has been widely used in various computer vision tasks, such as image classification [7, 57, 58],image segmentation [27, 67, 6], object detection [33, 32, 46], image generation [13, 23], video understanding [21, 24, 52], etc.
+Those remarkable achievements have demonstrated that Mamba can effectively address the challenges posed by CNNs and Transformers, including the local perception limitations of CNNs and the quadratic computational complexity of Transformers [55, 63].
+It benefits from a selective state space mechanism, which can efficiently capture global information and handle complex spatiotemporal dependencies with linear complexity [63].
+Specifically, Mamba has been conceptually summarized as being ideally suited for tasks involving long sequences and autoregression [60] due to its inherent RNN mechanism within the SSM [10, 9, 8].
+Inspired by Mamba’s capabilities for long sequence modeling and autoregression with linear complexity, we propose a novel optical flow estimation framework, MambaFlow, designed to address the challenges in CNNs or Transformer-based methods.
 As demonstrated in Fig.
-1, our method significantly reduces inference time while maintaining high accuracy.
-The framework introduces two core innovations: a feature enhancement Mamba (FEM) module, which optimizes feature representations to enhance global matching performance, and a flow propagation Mamba (FPM) module, which propagates flow information to tackle challenges in occluded regions.
-Specifically, our framework begins with a CNN backbone to extract robust features from the input image pair.
-We then propose the FEM module, which employs the Self-Mamba and Cross-Mamba mechanisms to strengthen the correlations between the extracted features, allowing for the capture of complex global information and long-range dependencies.
-Building on the enhanced features provided, a global matching module compute similarities and perform global optimization, to generate initial flow.
-This process fully utilizes global information, effectively tackling the challenges posed by large displacements.
-At last, we propose a Mamba-centric flow propagation method, the FPM module, which leverages feature similarity to guide propagation and produce the final accurate flow [36], tackling the occlusion issue in predicting optical flow.
+1, our method significantly reduces inference time while maintaining remarkblely high accuracy.The framework introduces two core components: (1) the PolyMamba module, which utilizes Self-Mamba, Cross-Mamba, and Multi-Layer Perceptron (MLP) to globally optimize feature representations with lightweight computation, and (2) the PulseMamba module, which effectively combines motion features with Mamba’s autoregressive ability to iteratively generate flow information, addressing challenges such as large displacement and occluded regions.
+Specifically, we begins with a CNN backbone to extract robust features from the input image pairs similar to traditional practice [49, 54, 20].
+Subsequently, we leverage the suggested PolyMamba to enhance the extracted features, allowing for the capture of complex global information and long-range dependencies.
+Building on the enhanced features, we further utilize a global matching mechanism [54, 64] to compute similarities for initial flow.
+At last, we employ PulseMamba to recurrently refine and yield the final precise flow [54].
+Our algorithm uses Mamba as the core components in both the encoding and decoding stage, making it an end-to-end optical flow estimation algorithm centered around Mamba.
+By leveraging Mamba’s capabilities in global modeling and autoregressive properties, as well as its computationally efficient characteristics, the proposed approach achieves a good trade-off between accuracy and performance.
 Our main contributions can be summarized as follows:
-  • We propose MambaFlow, a novel optical flow estimation architecture built on the Mamba framework.
-To the best of our knowledge, this is the first Mamba-centric architecture designed for end-to-end optical flow estimation.
-Our method achieves comparable accuracy to state-of-the-art algorithms while significantly reducing inference time.
-With a runtime of 0.113 seconds, our method is 18% faster than GMFlow and 10 times faster than MS-RAFT+.
-  • We introduce a feature enhancement Mamba (FEM) module, which integrates Self-Mamba and Cross-Mamba mechanisms to address insufficient feature representation.
-This module established a new Mamba-centric paradigm for accurate and efficient feature enhancement.
-  • We design a flow propagation Mamba (FPM) module for flow propagation that utilizes feature self-similarity to guide the final flow estimation.
-This module tackles the poor performance in recurrent refined methods by effectively propagating flow information, offering significant potential to enhance accuracy and speed, both currently and in the foreseeable future.
+  • We propose MambaFlow, a novel optical flow estimation architecture built on the Mamba framework both in feature enhancement and flow propagation.
+To our knowledge, this is the first Mamba-centric architecture designed for end-to-end optical flow estimation.
+Our method achieves accuracy comparable to state-of-the-art algorithms while significantly reducing training and inference time, which makes it more suitable for real-world application on resource-constrained devices.
+  • We introduce the PolyMamba module, which consists of a Self-Mamba, a Cross-Mamba, and a MLP in sequence, to globally enhance feature representation for both the reference image and its cross features with the matching image.
+As we will demonstrate, it effectively leverages the Mamba’s long sequence modeling capabilities while maintaining a linear computational load.
+  • We design the PulseMamba, which includes a motion encoder and a Mamba to facilitate flow propagation, addressing the computational overload, large displacement and occulusion issues in recurrently refined RNN-like methods.
+By leveraging the Mamba’s powerful autoregressive capability and linear complexity, it offers significant potential to enhance accuracy and speed simultaneously through ablation studies.
 2 Related Work
 2.1 Optical Flow Estimation
-The optical flow problem is traditionally approached as an energy minimization task.
-Based on the assumptions of brightness constancy and spatial smoothness, Horn and Schunck [8] introduced a variational method to tackle this problem.
-The following studies refined optical flow estimation by framing it as an optimization problem, improving accuracy and robustness through regularization that maximizes the visual similarity between corresponding pixels in the image.
+Optical flow estimation was traditionally formulated as an energy minimization task [1, 12, 38], with early variational methods like Horn and Schunck [12] relying on brightness constancy and spatial smoothness assumptions.
+Subsequent optimization-based approaches improved robustness through regularization [2].
 In the era of deep learning, following the introduction of FlowNet [5] as the first CNN-based optical flow method, numerous subsequent algorithms such as PWC-Net [44] and SpyNet [40], adopted a coarse-to-fine strategy for optical flow estimation.
-However, small displacements of fast movement were often missed via the inaccurate flow guidance at the coarse level.
-To overcome this obstacle, RAFT [31] proposed an iterative refinement method using full-field, achieving significant improvements.
-Constructing upon the innovation of RAFT, later works such as GMA [15] improved the handling of occluded regions by introducing a global motion aggregation module.
-Similarly, SeparableFlow [40] introduced a separable cost volume module that used a global motion aggregation layer to utilize global contextual cues and prior knowledge to resolve motion ambiguity in regions.
-Despite the advancements, CNN-based models, including those built upon the RAFT framework, still rely on multiple iterations of refinement to achieve high-precision optical flow estimation.
-This iterative process significantly increases computational overhead and slows inference speed, thereby limiting their practicality for efficiency-sensitive applications.
-In recent years, the Transformer architecture has demonstrated exceptional proficiency in modeling long-range dependencies and global information through its self-attention mechanism, making it a popular choice for optical flow estimation tasks.
-By incorporating more global and contextual information, it has significantly improved estimation accuracy.
-FlowFormer [10] further optimized model performance by encoding 4D cost tags into cost memory and applying alternating groups of Transformer layers in the latent space.
-CRAFT [29] strengthened the globality and semantic stability of features by leveraging semantic smoothing Transformer layers.
-GAFlow [22] added Gaussian attention to the existing Transformer block to emphasize local attributes during representation learning, highlight local neighborhoods containing fine-grained structural information, and strengthen motion affinity during matching.
-In addition, many models took advantage of the Transformer’s global modeling capabilities, adopting a global matching strategy to avoid multiple iterations of refinement, significantly enhancing efficiency.
-GMFlow [36] reformulated the optical flow problem as a global matching problem and obtained the optical flow using one-time global matching, which was more accurate and faster.
-Transflow [21] used spatial self-attention and cross-attention mechanisms between adjacent frames to effectively captured global dependencies, providing more accurate correlations and reliable matches in flow estimation.
-Still, despite these advances in accuracy, the Transformer-based optical flow methods face challenges due to the quadratic complexity of the attention mechanism led to huge computational complexity and memory requirements.
-This slows inference speed and restricts their use in resource-limited settings, creating practical bottlenecks.
-In our proposed MambaFlow architecture, we employ the Mamba module, which contains linear complexity and discriminative feature representation.
-This approach significantly reduces the overhead of computational resources and memory requirements while maintaining high accuracy comparable to Transformer-based methods, A remarkable breakthrough for the practical application of optical flow models.
-2.2 State Space Model in Optical Flow
+However, these approaches struggled with fast motions and small displacements due to inaccurate coarse-level guidance.
+RAFT [45] addressed this via iterative full-field refinement, inspiring variants such as GMA [20], SeparableFlow [62],and MS-RAFT [19], which introduced modules to better handle occlusion and motion ambiguity.
+Despite their accuracy, these CNN-based methods require heavy iterative computation and struggle with large displacements and complex motion, limiting their efficiency on resource-constrained devices.
+Transformers have emerged as powerful alternatives due to their ability to model global dependencies via self-attention.
+Models like FlowFormer [15], CRAFT [43], and GAFlow [31] leveraged Transformer architectures to enhance feature expressiveness and global matching.
+GMFlow [54] and TransFlow [30] further improved efficiency by reformulating optical flow as a one-shot global matching problem.
+However, the quadratic complexity of attention mechanisms results in high computational and memory costs, limiting deployment on lightweight systems.
+2.2 State Space Models
 State space models [11], which originated from control theory, have been effectively integrated with deep learning to capture long-range dependencies.
@@
 This superiority led to further exploration like Vision Mamba [68] and Vmamba [29], which applied SSMs to visual tasks such as image classification [29, 68], video understanding [25, 47], biomedical image segmentation [34, 53] and others [14, 17, 37, 39, 61].
-However, their applicability to optical flow estimation remained unclear until recently.
+However, their applicability to optical flow estimation has remained unexplored until recently.
 Lin et al.
-[18] pioneered Mamba into the scene flow estimation task, proposing an SSM-based iterative update module to replace the GRU module as a decoder.
-This approach can effectively integrate global motion information in the point cloud and significantly improved inference speed.
-Inspired by the Mamba’s linear complexity, we propose a Mamba-centric optical flow estimation architecture, aiming to replace the traditional attention mechanism and thus improve computational efficiency of the model.
-In this work, we explore the potential of Mamba for modeling sequential interactions in optical flow tasks and propose FEM module and FPM module.
-The FEM module enhances the features from pairs of images, capturing the dependency between them, while the FPM module refines optical flow predictions for occluded pixels by leveraging both the optical flow field and image structure similarity.
-By utilizing Mamba’s linear time complexity, our approach overcomes the quadratic complexity limitations of traditional Transformer models when processing long sequences, resulting in significantly improved inference efficiency.
+[26] pioneered the use of Mamba in the scene flow estimation task, proposing an SSM-based iterative update module to replace the GRU module as a decoder.
+Nevertheless, it was specifically designed for scene flow with point cloud feature and is not suitable for the generally used backbone optical flow methods.
+Inspired by the Mamba’s strength of long sequence modeling and autoregression with linear complexity, we propose a Mamba-centric architecture for optical flow estimation.
+Our goal is to design an end-to-end solution that utilizes Mamba for feature enhancement and flow optimization.
+During the feature enhancement phase, PolyMamba enhances features from pairs of images by sequentially employing Self-Mamba, Cross-Mamba, and MLP to capture global information and mutual dependencies.
+At the flow refinement phase, PulseMamba iteratively optimize the fina optical flow by leveraging a motion encoder and a SSM.
+These Mamba-centric designs guarantee the accuracy of our method while maintaining its high speed in training and inference, providing valuable insights for resource-constrained devices.
 Figure 2: The overall network architecture of our MambaFlow.
-First, we perform downsampling using convolutional layers and input the initially extracted features into the feature enhancement Mamba group (FEMG) module, which is composed of multiple feature enhancement Mamba (FEM) module.
-Each FEM module consists of Self-Mamba, Cross-Mamba, and MLP for feature enhancement.
-Next, global matching is applied to the output feature matrix for similarity matching, and the original optical flow is obtained through a softmax layer.
-Finally, the flow propagation Mamba (FPM) module is used to address inaccuracies in optical flow prediction caused by occlusion.
 3 Method
@@
 Simultaneously, the parallel scan algorithm enables Mamba to leverage the same parallel processing advantages as described in (5), thereby facilitating efficient training.
-Figure 3: The architecture of feature enhancement Mamba (FEM) module.
-It consists of three core components: Self-Mamba (a), which focuses on a single feature input, Cross-Mamba (b), which handles the interaction between two features, and MLP (c), which performs non-linear feature transformation.
 3.2 Overall Architecture
-Optical flow can be intuitively understood as a task aimed at finding pixel-wise correspondences between images.
+Optical flow can be intuitively understood as a task aimed at finding pixel-wise correspondences between consecutive images.
 As shown in Fig.
-2, our MambaFlow consists of three components: feature enhancement Mamba, global matching and flow propagation Mamba.
-Given two consecutive frames $\boldsymbol{I}_{1}$ and $\boldsymbol{I}_{2}$, we apply a shared-weight convolutional neural network to perform downsampling, obtaining dense feature maps $\boldsymbol{F}_{1},\boldsymbol{F}_{2}\in\mathbb{R}^{H\times W\times D}$.
-Here, $H,W,$ and $D$ denote the height, width, and feature dimension of the feature maps, respectively.
-Next, the image features are processed through the feature enhancement stage to obtain the enhanced features of the $l$-th layer, $\boldsymbol{\mathit{F}}_{q}$ and $\boldsymbol{\mathit{F}}_{v}\in\mathbb{R}^{H\times W\times D}$.
-This stage is composed of multiple stacked FEM modules, each consisting of Self-Mamba and Cross-Mamba blocks.
-Additionally, a Multi-Layer Perceptron (MLP) layer is introduced at the end of each group to further integrate local and global features and enhance feature representation.
-Finally, we employ the Global Matching method to predict optical flow and use our FPM module to significantly improve performance in occluded regions.
+2, our proposed framework, MambaFlow, is centered around two novel modules: the feature enhancement module PolyMamba and the flow refinement module PulseMamba.
+Given two input frames $\boldsymbol{I}_{1}$ and $\boldsymbol{I}_{2}$, we first extract dense feature maps $\boldsymbol{F}_{1},\boldsymbol{F}_{2}\in\mathbb{R}^{H\times W\times D}$ using a shared-weight convolutional encoder, where $H$, $W$, and $D$ denote the height, width, and feature dimension of the features.
+These features are then processed by the PolyMamba module, which stacks Self-Mamba and Cross-Mamba blocks to capture both intra-frame and inter-frame dependencies.
+To further enhance representation capacity, a lightweight MLP layer is appended after each block group for local-global feature fusion.
+The resulting features $\boldsymbol{\mathit{F}}_{q}$ and $\boldsymbol{\mathit{F}}_{v}\in\mathbb{R}^{H\times W\times D}$ are then used for global matching to estimate the initial flow.
+The initial prediction $\mathbf{V}_{\text{initial}}$ is then refined by the PulseMamba module, which incorporates cost volume and motion features into a stack of Mamba layers for iterative flow refinement.
+This module significantly improves performance in occluded or textureless regions, enhancing both the accuracy and robustness of optical flow estimation.
 3.3 Self-Mamba
 The input features $\boldsymbol{\mathit{F}}_{1},\boldsymbol{\mathit{F}}_{2}\in\mathbb{R}^{H\times W\times D}$ are enriched with learnable positional embeddings to retain spatial positional information.
-Subsequently, $\boldsymbol{\mathit{F}}_{1}$ and $\boldsymbol{\mathit{F}}_{2}$ are processed through a weight-sharing Self-Mamba block, generating $x_{q}$ and $x_{v}$, respectively.
-As shown in Fig.
-3 (a), the processing within each Self-Mamba block can be summarized as a bidirectional operation, where the forward and backward hidden states are computed and combined to enrich the feature representation.
+Subsequently, $\boldsymbol{\mathit{F}}_{1}$ and $\boldsymbol{\mathit{F}}_{2}$ are processed through a weight-sharing Self-Mamba block, yielding enhanced representations denoted as $\boldsymbol{\mathit{F}}_{1}^{\text{self}}$ and $\boldsymbol{\mathit{F}}_{2}^{\text{self}}$, respectively.
+The processing within each Self-Mamba block can be summarized as a bidirectional operation, where the forward and backward hidden states are computed and combined to enrich the feature representation.
 The overall process is expressed as:
@@
 […]
-Here, the Self-Mamba function represents the state space model used to compute the forward and backward hidden states, with the sequence $x$ being reversed to distinguish between the forward and backward directions.
+Here, the Self-Mamba function represents the state space model used to compute the forward and backward hidden states, with the sequence $\boldsymbol{\mathit{F}_{i}}$ being reversed to distinguish between the forward and backward directions.
 3.4 Cross-Mamba
-Mamba is specifically designed to capture long-term dependencies within a single sequence, but it cannot model the interrelations between two sets as effectively as the cross-attention mechanism in the Transformers.
-To overcome this limitation, we propose Cross-Mamba, an extension of the Mamba framework that adapts the cross-attention mechanism to better capture the interactions between two distinct sets.
-This idea, which builds upon the work by [33], is further extended in our approach.
-Since the parameters $\overline{\mathbf{A}}$, $\overline{\mathbf{B}}$, ${\mathbf{C}}$, and $\boldsymbol{\Delta}$ are associated with the input, $\overline{\mathbf{K}}$ in (5) is extended into an $M\times M$ matrix [1].
-The element $\alpha_{i,j}$, located at row $i$ and column $j$, is determined as follows:
-$\alpha_{i,j}=\mathbf{C_{i}}\prod_{k=j+1}^{i}\overline{\mathbf{A}}_{k}\overline{\mathbf{B}}_{j}.$ (11)
-Substituting the equations from (3) to (9), we can ultimately derive the output of Mamba as follows:
-$y_{i}=\sum_{j=1}^{i}\alpha_{i,j}x_{j}=\sum_{j=1}^{i}Q_{i}K_{j}H_{i,j}V_{j},$ (12)
-where
-$Q_{i}$ $=S_{C}(x_{i}),$ (13)
-$\begin{split}K_{j}=&\left(\text{Softplus}(S_{\Delta}(x_{j}))\mathbf{A}\right)^{-1}\\ &\left(\exp\left(\text{Softplus}(S_{\Delta}(x_{j}))\mathbf{A}\right)-\mathbf{I}\right)\\ &\left(\text{Softplus}(S_{\Delta}(x_{j}))S_{B}(x_{j})\right),\end{split}$ (14)
-$H_{i,j}$ $=\prod_{k=j+1}^{i}\exp\left(\text{Softplus}(S_{\Delta}(x_{k}))\mathbf{A}\right),$ (15)
-$V_{i}$ $=x_{i}.$ (16)
-Cross-attention is a crucial mechanism for modeling the relationships between two sets.
-It operates by computing interactions between the query ($Q$) from one set and the keys ($K$) and values ($V$) from another set, thereby identifying and focusing on the most relevant elements in the target set.By applying $x_{q}$ to (13), and $x_{v}$ to (14), (15) and (16), respectively, Cross-Mamba can be expressed in the following form:
-$y_{i}=\sum_{j=1}^{i}f_{m,q}(x_{q,i})f_{m,k}(x_{v,j})H_{i,j}x_{v,j}.$ (17)
-Cross-Mamba extends the original Mamba by expanding its application from a single input sequence to two input sequences, as shown in Fig. 3 (b). It is formally defined as follows:
-$y=\text{Cross-Mamba}(x_{q},x_{v}).$ (18)
-Similarly, we introduce the bidirectional mechanism of Self-Mamba into Cross-Mamba. The formulation is expressed as follows:
-[…]
-[…]
+While the Self-Mamba architecture is effective in modeling long-range dependencies within a single sequence, it lacks the capacity to capture inter-sequence relationships—an ability that is crucial in optical flow estimation, where information from two correlated feature maps must be jointly reasoned.
+To address this limitation, we introduce a cross-sequence interaction module based on selective state-space modeling, enabling rich bidirectional information exchange between paired features.
+Given the enhanced features $\boldsymbol{\mathit{F}}_{1}^{\text{self}},\boldsymbol{\mathit{F}}_{2}^{\text{self}}$, the main branch applies a depth-wise causal convolution and activation:
+$\boldsymbol{x}_{\mathrm{conv}}=\texttt{SiLU}\left(\texttt{DWConv}(\boldsymbol{\mathit{F}}_{1}^{\text{self}})\right),$ (11)
+while the modulation branch linearly projects the second stream:
+$\boldsymbol{x}_{\mathrm{mod}}=\texttt{Proj}(\boldsymbol{\mathit{F}}_{2}^{\text{self}}).$ (12)
+Their concatenation is projected into three parameter groups:
+$[\,\boldsymbol{\Delta},\;\mathbf{B},\;\mathbf{C}\,]=\texttt{Linear}([\boldsymbol{x}_{\mathrm{conv}};\;\boldsymbol{x}_{\mathrm{mod}}]),$ (13)
+The hidden state $\boldsymbol{y}$ is then updated through a SelectiveScan kernel, which dynamically integrates temporal and spatial information using parameterized state-space operations.
+A final linear projection yields the cross-attended representation:
+$\boldsymbol{\mathit{F}}^{\text{cross}}=\texttt{OutProj}(\boldsymbol{y}).$ (14)
+To model temporal symmetry, Cross-Mamba is applied in both forward and backward directions—just as in the Self-Mamba block—and the outputs are summed:
 […]
 3.5 Global Matching
-After extracting the enhanced features $\boldsymbol{\mathit{F}}_{q}$ and $\boldsymbol{\mathit{F}}_{v}$ from our FEM module, we construct a 4D cost volume of size $H\times W\times H\times W$ by computing the dot-product similarity between all pixel pairs from both features.
+After extracting the enhanced features $\boldsymbol{\mathit{F}}_{q}$ and $\boldsymbol{\mathit{F}}_{v}$ from our PolyMamba module, we construct a 4D cost volume of size $H\times W\times H\times W$ by computing the dot-product similarity between all pixel pairs from both features.
+The constructed 4D cost volume can be interpreted as a collection of $H\times W$ 2D cost maps, where each map evaluates the feature similarity between a single pixel in $\boldsymbol{\mathit{F}}_{q}$ and all pixels in $\boldsymbol{\mathit{F}}_{v}$.
+To obtain a matching distribution, a softmax operation is applied across the last two dimensions of the cost volume.
 The construction process is expressed as:
-$\mathbf{C}=\frac{\boldsymbol{\mathit{F}}_{q}\,\boldsymbol{\mathit{F}}_{v}^{\top}}{\sqrt{D}},$ (20)
+$\mathbf{M}=\text{softmax}(\mathbf{\frac{\boldsymbol{\mathit{F}}_{\mathit{q}}\,\boldsymbol{\mathit{F}}_{\mathit{v}}^{\top}}{\sqrt{D}}}).$ (16)
 where $D$ is the feature dimension.
-The constructed 4D cost volume can be interpreted as a collection of $H\times W$ 2D cost maps, where each map evaluates the feature similarity between a single pixel in $\boldsymbol{\mathit{F}}_{q}$ and all pixels in $\boldsymbol{\mathit{F}}_{v}$.
-To obtain a matching distribution, a softmax operation is applied across the last two dimensions of the cost volume:
-$\mathbf{M}=\text{softmax}(\mathbf{C}).$ (21)
-To estimate the optical flow, the correspondence $\hat{\mathbf{G}}$ is obtained by taking a weighted average of the 2D pixel grid coordinates $\mathbf{G}\in\mathbb{R}^{H\times W\times 2}$ in the target image with the matching distribution $M$:
-$\hat{\mathbf{G}}=\mathbf{M}\mathbf{G}.$ (22)
-Finally, the optical flow $\mathbf{V}_{\text{original}}$ is calculated as the difference between the corresponding pixel coordinates:
-$\mathbf{V}_{\text{original}}=\hat{\mathbf{G}}-\mathbf{G}.$ (23)
+To estimate the optical flow, the correspondence is obtained by taking a weighted average of the 2D pixel grid coordinates $\mathbf{G}\in\mathbb{R}^{H\times W\times 2}$ in the target image with the matching distribution $M$.
+Finally, the optical flow $\mathbf{V}_{\text{initial}}$ is calculated as the difference between the corresponding pixel coordinates:
+$\mathbf{V}_{\text{initial}}=\mathbf{M}\mathbf{G}-\mathbf{G}.$ (17)
 The global matching method effectively improves the accuracy and stability of optical flow estimation by capturing global contextual information.
-3.6 Flow Propagation Mamba
-To address the issue of occluded regions or pixels outside the image boundaries, we introduced a self-similarity-based module called the flow propagation Mamba (FPM) module (as illustrated in Fig.
-4), which leverages feature information from the source image to capture pixel-level similarity relationships.
-Since adjacent pixels typically exhibit similar features, they demonstrate higher similarity, thus providing a more robust solution for optical flow estimation.
-Specifically, by inputting $\boldsymbol{\mathit{F}}_{q}$ into the Mamba block, we enhance the features in occluded regions and further extract the weights of these occluded areas.
-This helps optimize the final optical flow estimation, thereby improving accuracy and robustness in complex scenarios.
-The formula is as follows:
-$\mathbf{V}_{\text{final}}=\text{softmax}\left(\text{Mamba}(\boldsymbol{\mathit{F}}_{q})\right)\mathbf{V}_{\text{initial}},$ (24)
-where $\mathbf{V}_{\text{initial}}\in\mathbb{R}^{H\times W\times 2}$ represents the optical flow predicted by the global matching.
-Figure 4: The architecture of flow propagation Mamba (FPM) module.
-Figure 5: Visual results on Sintel test set.
+3.6 PulseMamba
+To effectively address the challenges of occlusions and boundary conditions in optical flow estimation, we propose a novel module named PulseMamba.
+PulseMamba integrates cost volume information and motion features, achieving dynamic and context-aware optical flow propagation through iterative refinement.
+Given the initial optical flow $\mathbf{V}_{\text{initial}}$, the source image features $\boldsymbol{F}_{q}$, and the 4D cost volume (denoted as $\mathbf{C}$), PulseMamba proceeds as follows:
+First, motion features are extracted by encoding the initial flow and the 4D cost volume:
+$\boldsymbol{M}=\text{MotionEncoder}(\mathbf{V}_{\text{initial}},\mathbf{C}),$ (18)
+PulseMamba employs multiple stacked Mamba layers for iterative feature refinement.
+At each iteration $i$, the input is formed by concatenating the image features, motion features, and current flow prediction, and the flow is updated as:
+$\mathbf{V}^{(i)}=\text{Mamba}^{(i)}\left([\boldsymbol{F}_{q};\;\boldsymbol{M};\;\mathbf{V}^{(i-1)}]\right),$ (19)
+where $\mathbf{V}^{(0)}=\mathbf{V}_{\text{initial}}$.
+The output flow from each iteration becomes the input for subsequent refinement. After $N$ iterations, the final refined optical flow is obtained:
+$\mathbf{V}_{\text{final}}=\mathbf{V}^{(N)}.$ (20)
+Through this iterative and context-aware refinement, PulseMamba significantly improves the accuracy and robustness of optical flow estimation, particularly under complex conditions involving occluded regions and ambiguous boundaries.
+Table 1: Quantitative results on Sintel and KITTI.
+MambaFlow achieves at least 18% faster inference on the KITTI (test) dataset compared to other state-of-the-art methods with competitive performance.
+All compared methods were evaluated under the same training configuration, and a unified "C+T+S+K+H" training strategy was adopted to ensure the fairness of the results.
+For performance measurements, we used the normalized setting of batch size 1 and input resolution 540 × 960.
+In addition, all inference times are measured on the RTX 3090.
+The significant performance improvement highlights MambaFlow’s advantages in speed and efficiency.
+It is worth noting that while maintaining competitive accuracy on the Sintel and KITTI datasets, we have successfully achieved a significant improvement in speed, further verifying the superiority of MambaFlow in the speed-accuracy balance.
 Training Data Method Sintel KITTI Inference Cost
@@
 GMFlow [54] 1.74 2.90 9.32 4.7M 138.5ms
-MambaFlow (ours) 1.60 2.78 8.53 8.9M 113.5ms
-Table 1: MambaFlow achieves at least 18% faster inference on the KITTI (test) dataset compared to other state-of-the-art methods with competitive performance.
-All compared methods were evaluated under the same training configuration, and a unified ”C+T+S+K+H” training strategy was adopted to ensure the fairness of the results.
-For performance measurements, we used the normalized setting of batch size 1 and input resolution 540 × 960.
-In addition, all inference times are measured on the RTX 3090 graphics card.
-The significant performance improvement highlights MambaFlow’s advantages in speed and efficiency.
-It is worth noting that while maintaining competitive accuracy on the Sintel and KITTI datasets, we have successfully achieved a significant improvement in speed, further verifying the superiority of MambaFlow in the speed-accuracy balance.
+MambaFlow (ours) 1.43 2.45 8.06 11.2M 113.5ms
 4 Experiments
@@
 To gain deeper insights into the model’s performance under large motion magnitudes, we defined $s_{40+}$ as the EPE for pixels where the ground truth flow magnitude exceeds 40 pixels.
-setup Things (val) Sintel (train) Param (M)
-[…]
-full 6.45 - 2.23 - 3.41 - 8.1 -
-w/o self 6.89 $\uparrow$6.82.30 $\uparrow$3.13.49 $\uparrow$2.36.9 $\downarrow$14.8%
-w/o cross 11.10 $\uparrow$72.13.78 $\uparrow$69.55.16 $\uparrow$51.34.4 $\downarrow$45.7%
-w/o MLP 7.13 $\uparrow$10.52.41 $\uparrow$8.13.66 $\uparrow$7.35.2 $\downarrow$35.8%
-w/o pos 6.65 $\uparrow$3.12.29 $\uparrow$2.73.48 $\uparrow$2.18.1 -
-Table 2: FEM components. Ablation study on individual components, revealing that Cross-Mamba contributes the most.
-layers Sintel (train) KITTI (train) Param (M)
-[…]
-4 1.41 2.79 9.53 27.15 5.1
-6 1.37 2.72 9.16 25.38 6.9
-8 1.36 2.63 8.54 24.97 8.9
-10 1.35 2.68 8.51 24.67 10.8
-12 1.36 2.55 8.10 23.91 12.7
-24 1.34 2.50 7.96 22.97 23.4
-Table 3: Number of Mamba layers. Using 8 layers of Mamba strikes an optimal balance between accuracy and parameter efficiency.
 Implementation Details.
-Our model is constructed using the PyTorch library and trained on 4 NVIDIA GeForce RTX 3090 GPUs.
+Our model is constructed using the PyTorch library and trained on 8 NVIDIA GeForce RTX 3090 GPUs.
 Following the training strategy from previous works [54, 45], we use the AdamW optimizer.
@@
 Finally, we continue training on the KITTI-2015 training set for 200k iterations with the batch size and learning rate unchanged.
+Figure 3: Visual results on Sintel test set.
+Table 2: PolyMamba components. Ablation study on individual components, revealing that Cross-Mamba contributes the most.
+setup Things (val) Sintel (train) Param (M)
+[…]
+full 2.91 - 1.23 - 2.76 - 11.2 -
+w/o self 3.11 $\uparrow$6.91.27 $\uparrow$3.32.82 $\uparrow$2.28.9 $\downarrow$24.6%
+w/o cross 5.01 $\uparrow$72.22.08 $\uparrow$69.14.18 $\uparrow$51.45.4 $\downarrow$54.2%
+w/o MLP 3.22 $\uparrow$10.71.33 $\uparrow$8.12.96 $\uparrow$7.28.2 $\downarrow$30.5%
+w/o pos 3.00 $\uparrow$3.11.26 $\uparrow$2.42.82 $\uparrow$2.211.2 -
+Table 3: Number of PolyMamba blocks. Using 8 blocks of Mamba strikes an optimal balance between accuracy and parameter efficiency.
+blocks Sintel (train) KITTI (train) Param (M)
+[…]
+4 1.28 2.94 10.03 29.18 7.2
+6 1.24 2.87 9.64 27.23 9.2
+8 1.23 2.77 8.99 26.84 11.2
+10 1.22 2.82 8.96 26.52 13.2
+12 1.23 2.69 8.53 25.70 15.2
+24 1.21 2.63 8.38 24.69 27.3
 4.1 Comparison with State-of-the-Arts
@@
 Results on Sintel.
-MambaFlow achieved an EPE of 1.60 on the Sintel Clean dataset and an EPE of 2.78 on the Final dataset.
+MambaFlow achieved an EPE of 1.43 on the Sintel Clean dataset and an EPE of 2.45 on the Final dataset.
 Fig.
@@
 4.2 Ablation Analysis
-Components of FEM.
-We ablate different Fearture Enhancement Mamba components in Table 2.
+Components of PolyMamba.
+We ablate different PolyMamba components in Table 2.
 Among them, Cross-Mamba contributes the most as it models the mutual relationship between two features, a relationship that is otherwise missing in the features extracted from the convolutional backbone.
 Self-Mamba aggregates contextual cues within the same feature, resulting in an additional performance improvement, and positional information introduces location dependence into the matching process, helping to mitigate ambiguities present in feature-similarity-based matching alone.
-Removing the Multi-Layer Perceptron (MLP) reduces the number of parameters, but also results in a slight decrease in performance.
-Number of FEM layers.
-We conduct an ablation study on the number of FEM layers, as shown in Table 3.
-The experiment compares the impact of varying the number of FEM layers on overall performance.
-Among all the tested configurations, the FEMG module with 8 Mamba layers exhibits the best performance.
+Removing the MLP reduces the number of parameters, but also results in a slight decrease in performance.
+Number of PolyMamba blocks.
+We conduct an ablation study on the number of PolyMamba blocks, as shown in Table 3.
+The experiment compares the impact of varying the number of PolyMamba blocks on overall performance.
+Among all the tested configurations, the PolyMamba with 8 blocks achieves a good balance between accuracy and parameters.
+Number of PulseMamba iterations.
+As shown in Table 5, we investigated the impact of different iterations of PulseMamba on EPE on the Sintel dataset.
+As the number of iterations increases from 1 to 3, the overall EPE gradually decreases from 0.77 to 0.65; the error of the matched area decreases from 0.58 to 0.49, and the error of the unmatched area decreases significantly from 3.68 to 3.16, with the most obvious improvement.
+This trend shows that more iterations can improve the prediction accuracy of the occluded area.
+Weight sharing vs. non-sharing comparison.
+We compared Mamba modules with shared and independent weights at the same number of iterations.
+As shown in Table 5, the non-shared scheme achieved slightly better results on all metrics, matching and non-matching metrics.
+This is because the parameters of each layer can more flexibly adapt to feature changes between iterations.
+For example, in two iterations, shared weights reduced the model size by 7.5% (from 12.0 to 11.1), but caused the error to increase by 5.9% (from 0.68 to 0.72).
+These findings suggest that for complex matching tasks, it is reasonable to accept a moderate increase in the number of parameters due to the improvement in feature representation ability.
 Training iterations.
-The experimental results in Table 4 show that with the increase in the number of training iterations, our framework has achieved certain improvements in performance.
+As shown in Table 5, Mambaflow has achieved certain improvements in performance as the number of training iterations increases.
 This shows that more training iterations can effectively improve the performance of the model, making it more stable and accurate in subsequent tasks.
-training Things (val) Sintel (train) KITTI (train)
+Table 4: Ablations on PulseMamba designs.
+PulseMamba iterations share Mamba? Sintel (train, clean)
+all matched unmatched
+0 2.97 0.68 15.75
+1 0.77 0.58 3.68
 […]
-200K 3.04 1.19 2.60 9.02 25.92
-400K 2.81 1.14 2.57 8.76 25.18
-600K 2.73 1.10 2.52 8.14 24.18
-800K 2.61 1.09 2.56 7.57 23.51
-Table 4: Training iterations. MambaFlow benefits from more training iterations.
+[…]
+[…]
+[…]
+Table 5: Training iterations.
+[…]
 5 Conclusion
-In this paper, we explore for the first time the application of the advanced Mamba in optical flow estimation, and then propose a simple, accurate and efficient Mamba-centric optical flow method, MambaFlow, for end-to-end flow prediction.
-It consists of two components built on Mamba architecture: the FEM module and FPM module.
-The FEM is proposed for accurate and efficient feature enhancement, incorporating the Self-Mamba, Cross-Mamba mechanisms to tackle the issue of inadequate feature representation.
-Furthermore, the FPM that employs feature self-similarity to direct the ultimate flow is crafted for the flow propagation, addressing the challenge of traditional recurrent refined methods.
-Our method achieves comparable accuracy to state-of-the-art algorithms while significantly reducing inference time, with a runtime of 0.113 seconds—18% faster than GMFlow and 10 times faster than MS-RAFT+.
-We attribute this to the high accuracy and efficiency of Mamba in capturing features with local correlation while maintaining its global information.
-As far as we know, this is the first Mamba-centric architecture designed for end-to-end optical flow estimation.
-We are assured that additional variants of the Mamba-centric methods may emerge , possessing significant potential to boost both precision and velocity in the future.
+As Mamba continues to gain significant progresses in the field of computer vision, this paper explores for the first time to leverage it to estimate optical flow in an end-to-end paradigm.
+We suggest a Mamba-centric optical flow method, MambaFlow, which contributes two core Mamba-based components: PolyMamba and PulseMamba, proposed to accurately and efficiently enhance feature representation and direct the ultimate flow, respectively.The proposed approach achieves a good balance between accuracy and performance, comparing to the state-of-the-art methods.
+We attribute this to the Mamba’s capabilities in global modeling and autoregression, as well as its linear complexity, which offers greater potential for real-world deployment on resource-constrained devices, and provides valuable insights that may influence ongoing and future developments in this rapidly evolving field of optical flow.
+Limitations.
+While MambaFlow achieves results comparable to or superior to some state-of-the-art methods, there is still room for future improvement, as shown from the KITTI results in Table 1.
+Initially, it might not generalize well when there exists a substantial disparity between the training data and the test data.
+Secondly, it can still provide better accuracy under adverse conditions, such as motion blur, fog, and high-resolution, which remain unexplored.
+We will pretrain our model on large-scale datasets such as TartanAir [48] and VIPER [41], which can escalate its generalization ability.
+Furthermore, there are numerous outstanding variants of Mamba [63] have emerged recently, enabling us to adjust the entire architecture and its core Mamba-based components for more accurate estimations and to facilitate real-world applications.

### v3

@@
 To our knowledge, MambaFlow is the first architecture centered around the Mamba design tailored specifically for optical flow estimation.
-It comprises two key components: (1) PolyMamba, which optimizes feature representation; and (2) PulseMamba, which facilitates efficient flow information dissemination.
-Our extensive experiments demonstrate that MambaFlow achieves remarkable results.
-On the Sintel benchmark, MambaFlow records an endpoint error (EPE) of 1.43 and an inference speed of 0.113 seconds, surpassing the state-of-the-art methods including GMFlow (with 18.9% lower EPE and 18.1% faster inference), SeparableFlow (5% lower EPE and 50.5% faster), CRAFT (1.11% lower EPE and 76.5% faster), and DIP (0.7% lower EPE and 77.2% faster)—demonstrating stronger potential for real-world deployment on resource-constrained devices.
+It comprises two key components: (1) PolyMamba, which enhances feature representation through a dual-Mamba architecture, incorporating a Self-Mamba module for intra-token modeling and a Cross-Mamba module for inter-modality interaction, enabling both deep contextualization and effective feature fusion; and (2) PulseMamba, which leverages an Attention Guidance Aggregator (AGA) to adaptively integrate features with dynamically learned weights in contrast to naive concatenation, and then employs the intrinsic recurrent mechanism of Mamba to perform autoregressive flow decoding, facilitating efficient flow information dissemination.
+Extensive experiments demonstrate that MambaFlow achieves remarkable results comparable to mainstream methods on benchmark datasets.
+Compared to SEA-RAFT, MambaFlow attains higher accuracy on the Sintel benchmark, demonstrating stronger potential for real-world deployment on resource-constrained devices.
 The source code will be made publicly available upon acceptance of the paper.
-1 Introduction
+Introduction
 Optical flow estimation, a fundamental problem in computer vision, aims to compute the motion vectors of pixels between consecutive frames in a video sequence.
-It holds significant and widespread applications in areas such as action recognition [65, 56], video interpolation [16, 51, 28], autonomous driving [59], etc.
-Architectures based on Convolutional Neural Networks (CNNs) have achieved pioneering progress in the field of optical flow.
-FlowNet [5], as the first end-to-end CNNs-based flow network, revealed the immense potential of this learning paradigm.
-However, it faced limitations in handling large displacements and complex motions.
-Subsequently, PWC-Net [44] improved estimation accuracy in complex motion scenarios by incorporating pyramid feature extraction and a cost-volume construction, which effectively captured multi-scale local features.
-Nevertheless, it still encountered challenges in large displacement scenarios.
-To address these issues, RAFT [45] introduced a recurrent iterative refinement mechanism, which significantly enhanced the accuracy of optical flow estimation.
-However, such innovation brought a new challenge, which its optimization process relies on numerous iterations, leading to increased parameter counts and longer inference time.
-Many subsequent variants of the RAFT algorithm have attempted upgrates, such as MS-RAFT [19] and SEA-RAFT [49], but their ability to handle large displacements and complex motions still requires further improvement.
-To tackle the aforementioned issues, the Transformer architecture, known for its powerful global modeling capabilities, has been introduced into optical flow estimation tasks [15, 42, 30, 43].
-Most of them leverage attention mechanism to enhance feature representation, which capture long-range dependencies to significantly improve the accuracy.
-Notably, unlike optical flow algorithms based on RNNs for flow decoding [45, 20], GMFlow [54] utilizes the attention mechanism to iteratively decode the final optical flow, resulting in faster flow refinement.
-However, they encounter the quadratic computational complexity of Transformers both during the training and inference stages [54, 43, 15, 42, 30], which severely limits their efficiency in practical applications.
+It holds significant and widespread applications in areas such as action recognition (Zhao et al.
+2020; Xu et al.
+2019), video interpolation (Huang et al.
+2022b; Wu et al.
+2022; Liu, Liu, and Lin 2020), autonomous driving (Yi et al.
+2023), etc.
 Figure 1: End-point-error on Sintel (clean) vs.
 inference time (ms) and model params (M).
-All models were trained on FlyingChairs (C), FlyingThings3D (T), Sintel (S), KITTI (K), and HD1K (H) datasets, with testing performed using one NVIDIA GeForce RTX 3090 GPU.
 Compared to state-of-the-art models, our proposed model achieves comparable accuracy while significantly reducing inference time, demonstrating an optimal balance between computational efficiency and performance.
-Recently, Mamba [8, 10] has been widely used in various computer vision tasks, such as image classification [7, 57, 58],image segmentation [27, 67, 6], object detection [33, 32, 46], image generation [13, 23], video understanding [21, 24, 52], etc.
-Those remarkable achievements have demonstrated that Mamba can effectively address the challenges posed by CNNs and Transformers, including the local perception limitations of CNNs and the quadratic computational complexity of Transformers [55, 63].
-It benefits from a selective state space mechanism, which can efficiently capture global information and handle complex spatiotemporal dependencies with linear complexity [63].
-Specifically, Mamba has been conceptually summarized as being ideally suited for tasks involving long sequences and autoregression [60] due to its inherent RNN mechanism within the SSM [10, 9, 8].
+Architectures based on Convolutional Neural Networks (CNNs) have achieved pioneering progress in the field of optical flow.
+FlowNet (Dosovitskiy et al.
+2015), as the first end-to-end CNNs-based flow network, revealed the immense potential of this learning paradigm.
+However, it faced limitations in handling large displacements and complex motions.
+Subsequently, PWC-Net (Sun et al.
+2018) improved estimation accuracy in complex motion scenarios by incorporating pyramid feature extraction and a cost-volume construction, which effectively captured multi-scale local features.
+Nevertheless, it still encountered challenges in large displacement scenarios.
+To address these issues, RAFT (Teed and Deng 2020) introduced a recurrent iterative refinement mechanism, which significantly enhanced the accuracy of optical flow estimation.
+However, such innovation brought a new challenge, which its optimization process relies on numerous iterations, leading to increased parameter counts and longer inference time.
+Many subsequent variants of the RAFT algorithm have attempted upgrates (Jahedi et al.
+2024b; Wang, Lipson, and Deng 2025; Zheng et al.
+2022; Sun et al.
+2022; Dong, Cao, and Fu 2023; Jahedi et al.
+2024a), such as MS-RAFT and SEA-RAFT, but their ability to handle large displacements and complex motions still requires further improvement.
+To tackle the aforementioned issues, the Transformer architecture, known for its powerful global modeling capabilities, has been introduced into optical flow estimation tasks (Huang et al.
+2022a; Shi et al.
+2023; Lu et al.
+2023; Sui et al.
+2022).
+Most of them leverage attention mechanism to enhance feature representation, which capture long-range dependencies to significantly improve the accuracy.
+Notably, unlike optical flow algorithms based on RNNs for flow decoding (Teed and Deng 2020; Jiang et al.
+2021), GMFlow (Xu et al.
+2022) utilizes the attention mechanism to iteratively decode the final optical flow, resulting in faster flow refinement.
+However, they encounter the quadratic computational complexity of Transformers both during the training and inference stages (Xu et al.
+2022; Sui et al.
+2022; Huang et al.
+2022a; Shi et al.
+2023; Lu et al.
+2023), which severely limits their efficiency in practical applications.
+Recently, Mamba (Gu and Dao 2023; Gu et al.
+2021) has been widely used in various computer vision tasks, such as image classification (Ge et al.
+2024; Yang, Wang, and Chen 2024; Yang et al.
+2024),image segmentation (Liu et al.
+2024a; Zhou et al.
+2024; Du, Wang, and Chen 2024), object detection (Ma et al.
+2024; Ma, Zhang, and Shi 2024; Verma et al.
+2024), image generation (Hu et al.
+2024; Lambrechts et al.
+2024), video understanding (Jin et al.
+2024; Li et al.
+2024; Xiao et al.
+2024), etc.
+Those remarkable achievements have demonstrated that Mamba can effectively address the challenges posed by CNNs and Transformers, including the local perception limitations of CNNs and the quadratic computational complexity of Transformers (Xu et al.
+2024; Zhang et al.
+2024).
+It benefits from a selective state space mechanism, which can efficiently capture global information and handle complex spatiotemporal dependencies with linear complexity (Zhang et al.
+2024).
+Specifically, Mamba has been conceptually summarized as being ideally suited for tasks involving long sequences and autoregression (Yu and Wang 2024) due to its inherent RNN mechanism within the SSM (Gu et al.
+2021; Gu, Goel, and Ré 2021; Gu and Dao 2023).
 Inspired by Mamba’s capabilities for long sequence modeling and autoregression with linear complexity, we propose a novel optical flow estimation framework, MambaFlow, designed to address the challenges in CNNs or Transformer-based methods.
-As demonstrated in Fig.
-1, our method significantly reduces inference time while maintaining remarkblely high accuracy.The framework introduces two core components: (1) the PolyMamba module, which utilizes Self-Mamba, Cross-Mamba, and Multi-Layer Perceptron (MLP) to globally optimize feature representations with lightweight computation, and (2) the PulseMamba module, which effectively combines motion features with Mamba’s autoregressive ability to iteratively generate flow information, addressing challenges such as large displacement and occluded regions.
-Specifically, we begins with a CNN backbone to extract robust features from the input image pairs similar to traditional practice [49, 54, 20].
+As demonstrated in Figure 1, our method significantly reduces inference time while maintaining remarkblely high accuracy.The framework introduces two core components: (1) the PolyMamba module, which utilizes Self-Mamba, Cross-Mamba, and Multi-Layer Perceptron (MLP) to globally optimize feature representations with lightweight computation, and (2) the PulseMamba module, which first employs an Attention Guidance Aggregator (AGA) to adaptively integrate motion features using dynamically learned weights—instead of naive concatenation—and then leverages Mamba’s autoregressive capability to iteratively decode and refine optical flow, effectively addressing challenges such as large displacement and occluded regions.
+Specifically, we begins with a CNN backbone to extract robust features from the input image pairs similar to traditional practice (Wang, Lipson, and Deng 2025; Xu et al.
+2022; Jiang et al.
+2021).
 Subsequently, we leverage the suggested PolyMamba to enhance the extracted features, allowing for the capture of complex global information and long-range dependencies.
-Building on the enhanced features, we further utilize a global matching mechanism [54, 64] to compute similarities for initial flow.
-At last, we employ PulseMamba to recurrently refine and yield the final precise flow [54].
+Building on the enhanced features, we further utilize a global matching mechanism (Xu et al.
+2022; Zhao et al.
+2022) to compute similarities for initial flow.
+At last, we employ PulseMamba to recurrently refine and yield the final precise flow (Xu et al.
+2022).
 Our algorithm uses Mamba as the core components in both the encoding and decoding stage, making it an end-to-end optical flow estimation algorithm centered around Mamba.
@@
 As we will demonstrate, it effectively leverages the Mamba’s long sequence modeling capabilities while maintaining a linear computational load.
-  • We design the PulseMamba, which includes a motion encoder and a Mamba to facilitate flow propagation, addressing the computational overload, large displacement and occulusion issues in recurrently refined RNN-like methods.
+  • We design the PulseMamba, which integrates an Attention Guidance Aggregator for adaptive feature fusion and a Mamba-based autoregressive decoder, addressing the computational overload, large displacement and occulusion issues in recurrently refined RNN-like methods.
 By leveraging the Mamba’s powerful autoregressive capability and linear complexity, it offers significant potential to enhance accuracy and speed simultaneously through ablation studies.
-2 Related Work
-2.1 Optical Flow Estimation
-Optical flow estimation was traditionally formulated as an energy minimization task [1, 12, 38], with early variational methods like Horn and Schunck [12] relying on brightness constancy and spatial smoothness assumptions.
-Subsequent optimization-based approaches improved robustness through regularization [2].
-In the era of deep learning, following the introduction of FlowNet [5] as the first CNN-based optical flow method, numerous subsequent algorithms such as PWC-Net [44] and SpyNet [40], adopted a coarse-to-fine strategy for optical flow estimation.
+Related Work
+Optical Flow Estimation
+Optical flow estimation was traditionally formulated as an energy minimization task (Brox et al.
+2004; Horn and Schunck 1981; Papenberg et al.
+2006), with early variational methods like Horn and Schunck (Horn and Schunck 1981) relying on brightness constancy and spatial smoothness assumptions.
+Subsequent optimization-based approaches improved robustness through regularization (Bruhn, Weickert, and Schnörr 2005).
+In the era of deep learning, following the introduction of FlowNet (Dosovitskiy et al.
+2015) as the first CNN-based optical flow method, numerous subsequent algorithms such as PWC-Net (Sun et al.
+2018) and SpyNet (Ranjan and Black 2017), adopted a coarse-to-fine strategy for optical flow estimation.
 However, these approaches struggled with fast motions and small displacements due to inaccurate coarse-level guidance.
-RAFT [45] addressed this via iterative full-field refinement, inspiring variants such as GMA [20], SeparableFlow [62],and MS-RAFT [19], which introduced modules to better handle occlusion and motion ambiguity.
+RAFT (Teed and Deng 2020) addressed this via iterative full-field refinement, inspiring variants such as GMA (Jiang et al.
+2021), SeparableFlow (Zhang et al.
+2021), and MS-RAFT (Jahedi et al.
+2024b), which introduced modules to better handle occlusion and motion ambiguity.
 Despite their accuracy, these CNN-based methods require heavy iterative computation and struggle with large displacements and complex motion, limiting their efficiency on resource-constrained devices.
 Transformers have emerged as powerful alternatives due to their ability to model global dependencies via self-attention.
-Models like FlowFormer [15], CRAFT [43], and GAFlow [31] leveraged Transformer architectures to enhance feature expressiveness and global matching.
-GMFlow [54] and TransFlow [30] further improved efficiency by reformulating optical flow as a one-shot global matching problem.
+Models like FlowFormer (Huang et al.
+2022a), CRAFT (Sui et al.
+2022), and GAFlow (Luo et al.
+2023) leveraged Transformer architectures to enhance feature expressiveness and global matching.
+GMFlow (Xu et al.
+2022) and TransFlow (Lu et al.
+2023) further improved efficiency by reformulating optical flow as a one-shot global matching problem.
 However, the quadratic complexity of attention mechanisms results in high computational and memory costs, limiting deployment on lightweight systems.
-2.2 State Space Models
-State space models [11], which originated from control theory, have been effectively integrated with deep learning to capture long-range dependencies.
-Mamba [8] introduced an input adaptation mechanism to enhance the state space model, providing faster inference speed, higher throughput and overall indicators compared to the Transformers of the same size, especially in natural language.
-Related work on Mamba [50] has shown that it has comparable performance to the Transformer-based methods while significantly reducing computational complexity.
-This superiority led to further exploration like Vision Mamba [68] and Vmamba [29], which applied SSMs to visual tasks such as image classification [29, 68], video understanding [25, 47], biomedical image segmentation [34, 53] and others [14, 17, 37, 39, 61].
+State Space Models
+State space models (Gu, Goel, and Ré 2022), which originated from control theory, have been effectively integrated with deep learning to capture long-range dependencies.
+Mamba (Gu and Dao 2023) introduced an input adaptation mechanism to enhance the state space model, providing faster inference speed, higher throughput and overall indicators compared to the Transformers of the same size, especially in natural language.
+Related work on Mamba (Wu et al.
+2025) has shown that it has comparable performance to the Transformer-based methods while significantly reducing computational complexity.
+This superiority led to further exploration like Vision Mamba (Zhu et al.
+2024) and Vmamba (Liu et al.
+2024b), which applied SSMs to visual tasks such as image classification (Liu et al.
+2024b; Zhu et al.
+2024), video understanding (Li et al.
+2025; Wang et al.
+2023), biomedical image segmentation (Ma, Li, and Wang 2024; Xing et al.
+2024) and others (Hu et al.
+2025; Islam et al.
+2023; Nguyen et al.
+2022; Qin et al.
+2024; Zha et al.
+2024).
 However, their applicability to optical flow estimation has remained unexplored until recently.
 Lin et al.
-[26] pioneered the use of Mamba in the scene flow estimation task, proposing an SSM-based iterative update module to replace the GRU module as a decoder.
+(Lin et al.
+2024) pioneered the use of Mamba in the scene flow estimation task, proposing an SSM-based iterative update module to replace the GRU module as a decoder.
 Nevertheless, it was specifically designed for scene flow with point cloud feature and is not suitable for the generally used backbone optical flow methods.
@@
 Our goal is to design an end-to-end solution that utilizes Mamba for feature enhancement and flow optimization.
-During the feature enhancement phase, PolyMamba enhances features from pairs of images by sequentially employing Self-Mamba, Cross-Mamba, and MLP to capture global information and mutual dependencies.
-At the flow refinement phase, PulseMamba iteratively optimize the fina optical flow by leveraging a motion encoder and a SSM.
 These Mamba-centric designs guarantee the accuracy of our method while maintaining its high speed in training and inference, providing valuable insights for resource-constrained devices.
 Figure 2: The overall network architecture of our MambaFlow.
-3 Method
-3.1 Preliminaries
+Method
+Preliminaries
 SSMs are inspired by linear systems, mapping an input sequence $x_{t}\in\mathbb{R}$ to an output $y_{t}\in\mathbb{R}$ through a hidden state $h_{t}\in\mathbb{R}^{N}$.
@@
 Simultaneously, the parallel scan algorithm enables Mamba to leverage the same parallel processing advantages as described in (5), thereby facilitating efficient training.
-3.2 Overall Architecture
+Overall Architecture
 Optical flow can be intuitively understood as a task aimed at finding pixel-wise correspondences between consecutive images.
-As shown in Fig.
-2, our proposed framework, MambaFlow, is centered around two novel modules: the feature enhancement module PolyMamba and the flow refinement module PulseMamba.
+As shown in Figure 2, our proposed framework, MambaFlow, is centered around two novel modules: the feature enhancement module PolyMamba and the flow refinement module PulseMamba.
 Given two input frames $\boldsymbol{I}_{1}$ and $\boldsymbol{I}_{2}$, we first extract dense feature maps $\boldsymbol{F}_{1},\boldsymbol{F}_{2}\in\mathbb{R}^{H\times W\times D}$ using a shared-weight convolutional encoder, where $H$, $W$, and $D$ denote the height, width, and feature dimension of the features.
@@
 This module significantly improves performance in occluded or textureless regions, enhancing both the accuracy and robustness of optical flow estimation.
-3.3 Self-Mamba
+Self-Mamba
 The input features $\boldsymbol{\mathit{F}}_{1},\boldsymbol{\mathit{F}}_{2}\in\mathbb{R}^{H\times W\times D}$ are enriched with learnable positional embeddings to retain spatial positional information.
@@
 The overall process is expressed as:
-[…]
-[…]
+$\boldsymbol{\mathit{F}}_{i}^{\text{self}}$ $=\text{Self-Mamba}(\boldsymbol{\mathit{F}}_{i})$ (10)
+$=\text{Self-Forward}(\boldsymbol{\mathit{F}}_{i})+\text{Self-Backward}(\boldsymbol{\mathit{F}}_{i})$
+$\quad i\in\{1,2\}.$
 Here, the Self-Mamba function represents the state space model used to compute the forward and backward hidden states, with the sequence $\boldsymbol{\mathit{F}_{i}}$ being reversed to distinguish between the forward and backward directions.
-3.4 Cross-Mamba
+Cross-Mamba
 While the Self-Mamba architecture is effective in modeling long-range dependencies within a single sequence, it lacks the capacity to capture inter-sequence relationships—an ability that is crucial in optical flow estimation, where information from two correlated feature maps must be jointly reasoned.
@@
 Given the enhanced features $\boldsymbol{\mathit{F}}_{1}^{\text{self}},\boldsymbol{\mathit{F}}_{2}^{\text{self}}$, the main branch applies a depth-wise causal convolution and activation:
-$\boldsymbol{x}_{\mathrm{conv}}=\texttt{SiLU}\left(\texttt{DWConv}(\boldsymbol{\mathit{F}}_{1}^{\text{self}})\right),$ (11)
+$\boldsymbol{x}_{\mathrm{conv}}=\operatorname{SiLU}\left(\operatorname{DWConv}(\boldsymbol{\mathit{F}}_{1}^{\text{self}})\right),$ (11)
 while the modulation branch linearly projects the second stream:
-$\boldsymbol{x}_{\mathrm{mod}}=\texttt{Proj}(\boldsymbol{\mathit{F}}_{2}^{\text{self}}).$ (12)
+$\boldsymbol{x}_{\mathrm{mod}}=\operatorname{Proj}(\boldsymbol{\mathit{F}}_{2}^{\text{self}}).$ (12)
 Their concatenation is projected into three parameter groups:
-$[\,\boldsymbol{\Delta},\;\mathbf{B},\;\mathbf{C}\,]=\texttt{Linear}([\boldsymbol{x}_{\mathrm{conv}};\;\boldsymbol{x}_{\mathrm{mod}}]),$ (13)
+$\boldsymbol{\Delta},\;\mathbf{B},\;\mathbf{C}\,]=\operatorname{Linear}([\boldsymbol{x}_{\mathrm{conv}};\;\boldsymbol{x}_{\mathrm{mod}}]),$ (13)
 The hidden state $\boldsymbol{y}$ is then updated through a SelectiveScan kernel, which dynamically integrates temporal and spatial information using parameterized state-space operations.
 A final linear projection yields the cross-attended representation:
-$\boldsymbol{\mathit{F}}^{\text{cross}}=\texttt{OutProj}(\boldsymbol{y}).$ (14)
+$\boldsymbol{\mathit{F}}^{\text{cross}}=\operatorname{OutProj}(\boldsymbol{y}).$ (14)
 To model temporal symmetry, Cross-Mamba is applied in both forward and backward directions—just as in the Self-Mamba block—and the outputs are summed:
-[…]
-3.5 Global Matching
+$\text{Cross-Mamba}\left(\boldsymbol{\mathit{F}}_{1}^{\text{self}},\;\boldsymbol{\mathit{F}}_{2}^{\text{self}}\right)$ (15)
+$=\;\text{Forward}(\,\cdot\,)+\text{Backward}(\,\cdot\,).$
+Global Matching
 After extracting the enhanced features $\boldsymbol{\mathit{F}}_{q}$ and $\boldsymbol{\mathit{F}}_{v}$ from our PolyMamba module, we construct a 4D cost volume of size $H\times W\times H\times W$ by computing the dot-product similarity between all pixel pairs from both features.
@@
 The global matching method effectively improves the accuracy and stability of optical flow estimation by capturing global contextual information.
-3.6 PulseMamba
-To effectively address the challenges of occlusions and boundary conditions in optical flow estimation, we propose a novel module named PulseMamba.
-PulseMamba integrates cost volume information and motion features, achieving dynamic and context-aware optical flow propagation through iterative refinement.
-Given the initial optical flow $\mathbf{V}_{\text{initial}}$, the source image features $\boldsymbol{F}_{q}$, and the 4D cost volume (denoted as $\mathbf{C}$), PulseMamba proceeds as follows:
-First, motion features are extracted by encoding the initial flow and the 4D cost volume:
-$\boldsymbol{M}=\text{MotionEncoder}(\mathbf{V}_{\text{initial}},\mathbf{C}),$ (18)
-PulseMamba employs multiple stacked Mamba layers for iterative feature refinement.
-At each iteration $i$, the input is formed by concatenating the image features, motion features, and current flow prediction, and the flow is updated as:
-$\mathbf{V}^{(i)}=\text{Mamba}^{(i)}\left([\boldsymbol{F}_{q};\;\boldsymbol{M};\;\mathbf{V}^{(i-1)}]\right),$ (19)
-where $\mathbf{V}^{(0)}=\mathbf{V}_{\text{initial}}$.
-The output flow from each iteration becomes the input for subsequent refinement. After $N$ iterations, the final refined optical flow is obtained:
+Extra Data Method Sintel KITTI Inference Cost
+[…]
+RAFT ECCV’20 1.61 2.86 5.10 5.3M 140.7ms
+SeparableFlow ICCV’21 1.50 2.67 4.64 6.0M 229.1ms
+GMA ICCV’21 1.39 2.47 5.15 5.9M 183.3ms
+DIP CVPR’22 1.44 2.83 4.21 5.4M 498.9ms
+GMFlowNet CVPR’22 1.39 2.65 4.79 9.3M 244.3ms
+CRAFT CVPR’22 1.45 2.42 4.79 6.3M 483.4ms
+GMFlow CVPR’22 1.74 2.90 9.32 4.7M 138.5ms
+SKFlow NeurIPS’22 1.28 2.23 4.85 6.3M 331.9ms
+MatchFlow CVPR’23 1.33 2.64 4.72 14.8M 283.8ms
+MS-RAFT+ IJCV’24 1.23 2.68 4.15 16.0M 1108.2ms
+MambaFlow (ours) 1.20 2.62 7.34 20.5M 116.5ms
+CroCo-Pretrain CroCoFlow ICCV’23 1.09 2.44 3.64 437.4M 6422.0ms
+DDVM-Pretrain DDVM NeurIPS’23 1.75 2.48 3.26
+YouTube-VOS FlowFormer++ CVPR’23 1.07 1.94 4.52 16.1M 373.4ms
+VIPER CCMR+ WACV’24 1.07 2.10 3.86 11.5M OOM
+TartanAir SEA-RAFT(M) ECCV’24 1.44 2.86 4.64 18.8M 71.0ms
+TartanAir SEA-RAFT(L) ECCV’24 1.31 2.60 4.30 18.8M 108.0ms
+Table 1: Quantitative results on Sintel and KITTI benchmarks.
+All methods are trained on the unified C+T+S+K+H datasets as the baseline.
+The ”Extra Data” column indicates additional pretraining datasets.
+Evaluation is conducted with batch size 1, input resolution $540\times 960$, and an RTX 3090 GPU.
+Figure 3: Visual results on Sintel test set.
+PulseMamba
+To effectively address occlusions and ambiguous boundary regions in optical flow estimation, we propose PulseMamba, an autoregressive refinement module featuring an Attention Guidance Aggregator (AGA) for adaptive feature integration.
+While conventional feature concatenation equally weights all spatial information, it often neglects dynamically relevant regions.
+To address this limitation, AGA adaptively assigns spatial attention weights to each feature, enabling context-aware flow estimation.
+Given initial optical flow $\mathbf{V}_{\text{initial}}$, source image features $\boldsymbol{F}_{q}$, and a 4D correlation cost volume $\mathbf{C}$, PulseMamba first extracts motion features:
+$\boldsymbol{M}=\text{MotionEncoder}(\mathbf{V}_{\text{initial}},\mathbf{C}).$ (18)
+AGA adaptively integrates motion features $\boldsymbol{M}$, context features $\boldsymbol{F}_{q}$, and hidden-state features $\boldsymbol{h}$ into a contextually informative representation.
+Specifically, AGA first aligns these input features to a unified embedding dimension through individual convolutional projections.
+Then, concatenated aligned features are processed via convolutional and GELU layers to produce spatial attention maps $\boldsymbol{A}$:
+$\boldsymbol{M};\,\boldsymbol{F}_{q};\,\boldsymbol{h}]\xrightarrow{\mathrm{Conv}_{1}}\xrightarrow{\mathrm{GELU}}\xrightarrow{\mathrm{Conv}_{2}}\xrightarrow{\mathrm{Softmax}}\boldsymbol{A.}$ (19)
+Finally, the aggregated feature $\boldsymbol{x}_{\text{AGA}}$ is obtained by a weighted sum:
+$\boldsymbol{x}_{\text{AGA}}=\sum_{f\in\{M,F_{q},h\}}\boldsymbol{A}_{f}\odot\boldsymbol{f},$ (20)
+where $\boldsymbol{A}_{f}$ represents spatial attention weights corresponding to each feature type.
+The aggregated feature $\boldsymbol{x}_{\text{AGA}}$ is then iteratively refined by PulseMamba using a Mamba layer to update the hidden-state representation:
+$\boldsymbol{h}^{(i)}=\text{Mamba}(\boldsymbol{x}_{\text{AGA}}^{(i-1)}).$ (21)
+Subsequently, the flow increment $\Delta\mathbf{V}^{(i)}$ is predicted, and the optical flow is iteratively updated:
+$\Delta\mathbf{V}^{(i)}$ $=\text{FlowHead}(\boldsymbol{h}^{(i)})$ (22)
+$\mathbf{V}^{(i)}$ $=\mathbf{V}^{(i-1)}+\Delta\mathbf{V}^{(i)}.$
+After NN iterations, PulseMamba produces the final refined optical flow:
 $\mathbf{V}_{\text{final}}=\mathbf{V}^{(N)}.$ (23)
-Through this iterative and context-aware refinement, PulseMamba significantly improves the accuracy and robustness of optical flow estimation, particularly under complex conditions involving occluded regions and ambiguous boundaries.
-Table 1: Quantitative results on Sintel and KITTI.
-MambaFlow achieves at least 18% faster inference on the KITTI (test) dataset compared to other state-of-the-art methods with competitive performance.
-All compared methods were evaluated under the same training configuration, and a unified "C+T+S+K+H" training strategy was adopted to ensure the fairness of the results.
-For performance measurements, we used the normalized setting of batch size 1 and input resolution 540 × 960.
-In addition, all inference times are measured on the RTX 3090.
-The significant performance improvement highlights MambaFlow’s advantages in speed and efficiency.
-It is worth noting that while maintaining competitive accuracy on the Sintel and KITTI datasets, we have successfully achieved a significant improvement in speed, further verifying the superiority of MambaFlow in the speed-accuracy balance.
-Training Data Method Sintel KITTI Inference Cost
-[…]
-C+T+S+K+H RAFT [45] 1.61 2.86 5.10 5.3M 140.7ms
-SeparableFlow [62] 1.50 2.67 4.64 6.0M 229.1ms
-GMA [20] 1.39 2.47 5.15 5.9M 183.3ms
-DIP [66] 1.44 2.83 4.21 5.4M 498.9ms
-GMFlowNet [64] 1.39 2.65 4.79 9.3M 244.3ms
-CRAFT [43] 1.45 2.42 7.79 6.3M 483.4ms
-MS-RAFT [18] 1.37 2.67 4.88 13.5M 627.3ms
-MatchFlow [4] 1.33 2.64 4.72 14.8M 283.8ms
-MS-RAFT+ [19] 1.23 2.68 4.15 16.0M 1108.2ms
-GMFlow [54] 1.74 2.90 9.32 4.7M 138.5ms
-MambaFlow (ours) 1.43 2.45 8.06 11.2M 113.5ms
-4 Experiments
+Through dynamic attention-guided feature integration and autoregressive refinement, PulseMamba effectively enhances the robustness and precision of optical flow estimation, especially under challenging scenarios.
+Experiments
 Datasets.
-Following the previous methodology, we first pre-train on the FlyingChairs (Chairs)[5] and FlyingThings3D (Things)[35] datasets.
-For fine-tuning, we adopt a mixed training set consisting of KITTI [36], HD1K [22], FlyingThings3D, and Sintel [3] training sets.
+Following the previous methodology, we first pre-train on the FlyingChairs (Chairs)(Dosovitskiy et al.
+2015) and FlyingThings3D (Things)(Mayer et al.
+2016) datasets.
+For fine-tuning, we adopt a mixed training set consisting of KITTI (Menze and Geiger 2015), HD1K (Kondermann et al.
+2016), FlyingThings3D, and Sintel (Butler et al.
+2012) training sets.
 Finally, we further fine-tune the model on the KITTI training set and report the performance on the respective online benchmarks.
@@
 Our model is constructed using the PyTorch library and trained on 8 NVIDIA GeForce RTX 3090 GPUs.
-Following the training strategy from previous works [54, 45], we use the AdamW optimizer.
+Following the training strategy from previous works (Xu et al.
+2022; Teed and Deng 2020), we use the AdamW optimizer.
 The model is first pre-trained on the Chairs dataset for 100k iterations with a batch size of 16 and a learning rate of $4\times 10^{-4}$.
@@
 Finally, we continue training on the KITTI-2015 training set for 200k iterations with the batch size and learning rate unchanged.
-Figure 3: Visual results on Sintel test set.
-Table 2: PolyMamba components. Ablation study on individual components, revealing that Cross-Mamba contributes the most.
-setup Things (val) Sintel (train) Param (M)
+Comparison with State-of-the-Arts
+We compared the performance of different optical flow algorithms on the Sintel and KITTI datasets, and the experimental results are shown in Table 1.
 […]
-full 2.91 - 1.23 - 2.76 - 11.2 -
-w/o self 3.11 $\uparrow$6.91.27 $\uparrow$3.32.82 $\uparrow$2.28.9 $\downarrow$24.6%
-w/o cross 5.01 $\uparrow$72.22.08 $\uparrow$69.14.18 $\uparrow$51.45.4 $\downarrow$54.2%
-w/o MLP 3.22 $\uparrow$10.71.33 $\uparrow$8.12.96 $\uparrow$7.28.2 $\downarrow$30.5%
-w/o pos 3.00 $\uparrow$3.11.26 $\uparrow$2.42.82 $\uparrow$2.211.2 -
-Table 3: Number of PolyMamba blocks. Using 8 blocks of Mamba strikes an optimal balance between accuracy and parameter efficiency.
+Among the methods that utilize extra pretraining data, SEA-RAFT is recognized as one of the fastest baselines.
+Our model achieves a similar inference speed to SEA-RAFT(L).
+Despite the use of additional TartanAir data in SEA-RAFT(L), MambaFlow achieves a lower EPE on Sintel Clean, with an 8.4% reduction compared to SEA-RAFT(L).
+For other methods that utilize large-scale extra pretraining data, such as FlowFormer++ (YouTube-VOS(Xu et al.
+2018)), CroCoFlow(Weinzaepfel et al.
+2023) (CroCo-Pretrain), and others, MambaFlow achieves comparable accuracy while maintaining significantly faster inference.
+Figure 3 presents a qualitative analysis of optical flow estimation, comparing MambaFlow with other state-of-the-art approaches on the Sintel test set.
+Overall, these results demonstrate that MambaFlow achieves a favorable trade-off between accuracy and efficiency, and delivers performance comparable to state-of-the-art methods, regardless of whether additional pretraining data is utilized.
+Setup Things (val) Sintel (train) Param
+clean clean final (M)
+full 2.22 1.15 3.05 20.5
+w/o self 2.40 1.24 3.12 18.2
+w/o cross 3.84 2.00 4.61 14.7
+w/o MLP 2.46 1.25 3.16 17.5
+w/o pos 2.29 1.18 3.11 20.5
+Table 2: Ablations on the Components of PolyMamba
 Blocks Sintel (train) KITTI (train) Param (M)
 […]
-4 1.28 2.94 10.03 29.18 7.2
-6 1.24 2.87 9.64 27.23 9.2
-8 1.23 2.77 8.99 26.84 11.2
-10 1.22 2.82 8.96 26.52 13.2
-12 1.23 2.69 8.53 25.70 15.2
-24 1.21 2.63 8.38 24.69 27.3
-4.1 Comparison with State-of-the-Arts
-We compared the performance of different optical flow algorithms on the Sintel and KITTI datasets, and the experimental results are shown in Table 1.
-Results on Sintel.
-MambaFlow achieved an EPE of 1.43 on the Sintel Clean dataset and an EPE of 2.45 on the Final dataset.
-Fig.
-3 presents a qualitative analysis of optical flow estimation, comparing MambaFlow with other state-of-the-art approaches on the Sintel test set.
-The results show that MambaFlow’s performance on Clean and Final is at the same level as current mainstream methods such as RAFT and GMA, demonstrating its strong adaptability in both simple and complex scenarios.
-Results on KITTI.
-MambaFlow achieves an F1-all score of 8.53% on the KITTI dataset while demonstrating a significant advantage in inference speed.
-Specifically, MambaFlow requires only 113.5ms for inference on an RTX 3090, making it 18.1% faster than GMFlow (138.5ms), which is the fastest among the compared methods.
-Furthermore, MambaFlow runs 2.5 times faster than MatchFlow (283.8ms) and nearly 10 times faster than MS-RAFT+ (1108.2ms), while maintaining competitive accuracy.
-This demonstrates that MambaFlow is able to maintain comparable accuracy to mainstream methods while significantly reducing inference time when handling larger and more challenging scenarios.
-4.2 Ablation Analysis
+4 1.20 3.21 7.54 22.28 16.5
+6 1.16 3.17 7.27 20.81 18.5
+8 1.15 3.05 6.76 20.50 20.5
+10 1.14 3.11 6.71 20.61 22.5
+12 1.15 2.96 6.45 18.90 24.5
+Table 3: Ablations on the Number of PolyMamba Blocks
+PulseMamba
+iterations
+use
+AGA?
+Sintel (train, clean)
+all matched unmatched
+0 1.96 0.68 15.75
+1 0.62 0.38 3.22
+0.72 0.45 3.58
+2 0.56 0.34 2.99
+0.60 0.36 3.16
+3 0.55 0.33 2.93
+0.57 0.35 2.91
+Table 4: Ablations on PulseMamba Designs
+Ablation Analysis
 Components of PolyMamba.
@@
 Among all the tested configurations, the PolyMamba with 8 blocks achieves a good balance between accuracy and parameters.
-Number of PulseMamba iterations.
-As shown in Table 5, we investigated the impact of different iterations of PulseMamba on EPE on the Sintel dataset.
-As the number of iterations increases from 1 to 3, the overall EPE gradually decreases from 0.77 to 0.65; the error of the matched area decreases from 0.58 to 0.49, and the error of the unmatched area decreases significantly from 3.68 to 3.16, with the most obvious improvement.
-This trend shows that more iterations can improve the prediction accuracy of the occluded area.
-Weight sharing vs. non-sharing comparison.
-We compared Mamba modules with shared and independent weights at the same number of iterations.
-As shown in Table 5, the non-shared scheme achieved slightly better results on all metrics, matching and non-matching metrics.
-This is because the parameters of each layer can more flexibly adapt to feature changes between iterations.
-For example, in two iterations, shared weights reduced the model size by 7.5% (from 12.0 to 11.1), but caused the error to increase by 5.9% (from 0.68 to 0.72).
-These findings suggest that for complex matching tasks, it is reasonable to accept a moderate increase in the number of parameters due to the improvement in feature representation ability.
-Training iterations.
-As shown in Table 5, Mambaflow has achieved certain improvements in performance as the number of training iterations increases.
-This shows that more training iterations can effectively improve the performance of the model, making it more stable and accurate in subsequent tasks.
-Table 4: Ablations on PulseMamba designs.
-PulseMamba iterations share Mamba? Sintel (train, clean)
-all matched unmatched
-0 2.97 0.68 15.75
-1 0.77 0.58 3.68
-[…]
-[…]
-[…]
-[…]
-Table 5: Training iterations.
-[…]
-5 Conclusion
+Ablation Study on PulseMamba and AGA.
+Table 4 shows that applying PulseMamba iterations brings substantial gains, particularly for unmatched pixels.
+A single iteration reduces overall EPE from 1.96 to 0.62, confirming the effectiveness of the design, and additional iterations further reduce the overall error.
+AGA consistently outperforms simple concatenation across all configurations; for example, with two iterations, the overall EPE decreases from 0.60 (concatenation) to 0.56 (AGA), demonstrating the benefit of adaptive feature fusion.
+Based on these results, we adopt two PulseMamba iterations with AGA as the default in our final model.
+Conclusion
 As Mamba continues to gain significant progresses in the field of computer vision, this paper explores for the first time to leverage it to estimate optical flow in an end-to-end paradigm.
-We suggest a Mamba-centric optical flow method, MambaFlow, which contributes two core Mamba-based components: PolyMamba and PulseMamba, proposed to accurately and efficiently enhance feature representation and direct the ultimate flow, respectively.The proposed approach achieves a good balance between accuracy and performance, comparing to the state-of-the-art methods.
+We suggest a Mamba-centric optical flow method, MambaFlow, which achieves a good balance between accuracy and performance compared to state-of-the-art methods.
 We attribute this to the Mamba’s capabilities in global modeling and autoregression, as well as its linear complexity, which offers greater potential for real-world deployment on resource-constrained devices, and provides valuable insights that may influence ongoing and future developments in this rapidly evolving field of optical flow.
 Limitations.
-While MambaFlow achieves results comparable to or superior to some state-of-the-art methods, there is still room for future improvement, as shown from the KITTI results in Table 1.
-Initially, it might not generalize well when there exists a substantial disparity between the training data and the test data.
-Secondly, it can still provide better accuracy under adverse conditions, such as motion blur, fog, and high-resolution, which remain unexplored.
-We will pretrain our model on large-scale datasets such as TartanAir [48] and VIPER [41], which can escalate its generalization ability.
-Furthermore, there are numerous outstanding variants of Mamba [63] have emerged recently, enabling us to adjust the entire architecture and its core Mamba-based components for more accurate estimations and to facilitate real-world applications.
+While MambaFlow achieves results comparable to some state-of-the-art methods, there is still room for improvement, as indicated by the KITTI results in Table 1.
+The model’s generalization may be limited when there is a substantial gap between training and test data, and its performance under challenging conditions such as motion blur, fog, or high-resolution images remains unexplored.
+We plan to pretrain on large-scale datasets like TartanAir (Wang et al.
+) and VIPER (Richter, Hayder, and Koltun 2017) to address these issues.
+Additionally, with the emergence of many Mamba variants (Zhang et al.
+2024), we see potential to further refine the architecture and core Mamba-based components for improved accuracy and real-world applicability.